### v1

# Prompting Techniques for Reducing Social Bias in LLMs through System 1 and System 2 Cognitive Processes

Abstract
Dual process theory posits that human cognition arises via two systems. System 1, which is a quick, emotional, and intuitive process, which is subject to cognitive biases, and System 2, a slow, onerous, and deliberate process. NLP researchers often compare zero-shot prompting in LLMs to System 1 reasoning and chain-of-thought (CoT) prompting to System 2. In line with this interpretation, prior research has found that using CoT prompting in LLMs leads to reduced gender bias. We investigate the relationship between bias, CoT prompting, and dual process theory in LLMs directly. We compare zero-shot, CoT, and a variety of dual process theory-based prompting strategies on two bias datasets spanning nine different social bias categories. We also use human and machine personas to determine whether the effects of dual process theory in LLMs are based on modeling human cognition or inherent to the system. We find that a human persona, System 2, and CoT prompting all tend to reduce social biases in LLMs, though the best combination of features depends on the exact model and bias category—resulting in up to a 13 percent drop in stereotypical judgments by an LLM.111 Datasets and code are available at https://github.com/kamruzzaman15/Prompting-Techniques-to-Reduce-Social-Bias-in-LLMs
1 Introduction
Figure 1: Example of Standard Prompting and Human Persona with System 2 Prompting for Llama 2 model in the race bias category
In recent years, large language models (LLMs) like GPT-4 (Achiam et al., 2023), ChatGPT (Brown et al., 2020), Llama 2 (Touvron et al., 2023) have revolutionized many aspects of technology and society. These models display remarkable linguistic capabilities, crafting responses that not only mimic human language but also exhibit a depth of understanding previously unattainable in machine learning (Karanjai and Shi, 2024). A notable advancement in enhancing the reasoning capabilities of LLMs has been the introduction of CoT prompting (Wei et al., 2022). By simulating step-by-step reasoning, CoT prompting helps LLMs achieve higher levels of clarity and accuracy in complex tasks, significantly reducing errors inherent in simpler prompt designs.
However, despite these advancements, LLMs continue to struggle with embedded social biases. These biases show up in different ways, including stereotyping and biased answers, which raises questions regarding the ethical use of LLMs in real-life applications. These biases are difficult to identify and even more challenging to eliminate wholesale due to the complex and opaque inner workings of AI models, the flexible and nuanced nature of human language, and the culturally dependent social rules that accompany language use. This task of mitigating social biases in LLMs is paramount to ensuring fairness and inclusivity in AI-driven communication and decisions. Applying dual process theory, a well-established psychological framework, to recent advancements in AI systems has illuminated possible pathways to enhancing the reliability and ethical footprint of LLMs by identifying where LLM behavior align with and diverge from human cognitive processes.
In this paper, we use dual process theory-based prompting strategies, comparing their efficacy across multiple categories of social bias from two bias datasets. Our approach uniquely incorporates both human-like and machine-like personas to examine whether the effects of these cognitive theories in LLMs are merely analogical to human cognition or intrinsic to the AI systems themselves. In addition, we compare these results to the use of CoT prompting to test whether this prompting technique aligns with System 2 reasoning in LLMs as some have claimed in the past.
Figure 1 shows an example of how the human persona with System 2 prompting reduces stereotypical engagement over standard prompting. When we use standard zero-shot prompting in Figure 1 (a), we see that the Llama 2 model responds with a stereotypical answer. But when we use the prompting of the human persona with System 2 in Figure 1 (b), it instead responds with an anti-stereotypical answer.
This paper’s contributions are the following.
  • We explore the effects of 12 different prompting techniques including CoT, System 1, System 2, and Persona, across nine distinct social bias categories (ageism, beauty, beauty with profession, gender, institutional, nationality, profession, race, religion) in 5 cutting-edge LLMs.
  • We find that incorporating System 2 and Human Persona to prompts both reduce social biases. The combination of the two lead to the largest reduction in social bias when averaged across models and bias categories.
  • In contradiction to the assumptions made by researchers in the past (Hagendorff et al., 2023), we find that CoT prompting does not behave similarly to prompts that directly model System 2 for social biases. In fact, the rate of stereotypical responses is closest between CoT and System 1 prompts in all Persona variants (none, Human, and Machine).
2 Related Work
Recent studies have explored how reasoning in LLMs can exhibit biases similar to human cognitive processes. Hagendorff et al. (2023) look into reasoning biases that are similar to humans in LLMs. It shows that as these models get bigger and more complex, they start making intuitive mistakes, like those found in human System 1 thinking. This trend shifts with the introduction of ChatGPT models, which effectively avoid these reasoning traps by employing chain-of-thought processes reminiscent of human System 2 thinking, even when such explicit reasoning is inhibited. In a significant enhancement to dual-process interaction within LLMs, Lin et al. (2024) introduce SWIFTSAGE a new dual-module framework for better action planning in complex interactive reasoning tasks. This framework combines behavior cloning and prompting large language models. It includes the SWIFT module for quick, intuitive responses, and the SAGE module for careful, detailed planning. Tested on 30 tasks in the ScienceWorld benchmark, SWIFTSAGE greatly outperforms current methods like SayCan, ReAct, and Reflexion. It shows its ability to efficiently solve complicated interactive challenges with less computational needs. Furthermore, Wei et al. (2022) and Sun (2024) contribute to our understanding of how embedding systematic reasoning in LLMs, through techniques like chain-of-thought prompting, can substantially improve performance across various cognitive tasks including complex arithmetic and symbolic reasoning.
Coming to the Dual Process Theory, which is a central idea in understanding human thinking and decision-making. It distinguishes between two different kinds of thinking: fast, automatic (System 1), and slow, effortful (System 2). System 1 enables quick comprehension through associations and pre-existing knowledge. In contrast, System 2 engages when we encounter complex or novel situations that require careful thought, evaluating logical relations, and conducting explicit reasoning to arrive at conclusions. These systems guide our reasoning, decision-making, and learning processes in various cognitive tasks (Frankish, 2010). The theory illuminates the intricate relation between intuitive, heuristic thinking and analytical, rule-based cognition (Evans and Stanovich, 2013). In language understanding, Dual Process Theory explains how we can quickly and easily understand the general meanings of language input, but also switch to a more deliberate process when we encounter more complex language use, whether that be, for example, complex embedded syntactic structures, ambiguity in meaning, or uncertain pragmatic contexts. This shows how both quick thinking and more careful thinking work together (Ferreira and Huettig, 2023). Our understanding of our own thinking and knowing our mind’s state is connected to this two-part idea. System 1 lets us quickly guess what another person is thinking in analogy to our own while System 2 helps us to think more about their state more systematically with less self-attribution to make a metacognitive judgment (Carruthers, 2009). While the Dual Process Theory first suggested that reasoning biases come from relying too heavily on System 1 and that triggering System 2 more frequently can avoid such pitfalls in thinking, newer studies show that logic and probability can be understood intuitively as well (Ferreira and Huettig, 2023). Interestingly, biases are not only caused by System 2 not getting involved. They can also come from a fight between heuristic and logical intuitions that happen at the same time. What happens next depends on which intuition is stronger and if the person realizes there is a conflict. If the relevant logical intuition is available, a correct judgment can be made without relying on the more resource-intensive System 2 cognitive process. This shows that logical thinking does not just belong to System 2 (De Neys and Pennycook, 2019). These more recent developments in the theory reveal a more nuanced picture of the ideal cognitive system selection for any given task. Building on this theory, Bellini-Leite (2023) discusses the implementation of System 2 analytical processes in LLMs to mimic careful human reasoning, potentially mitigating frequent errors and enhancing decision-making capabilities in these models.
Recent research on LLMs has found that assigning personas to LLMs can notably impact their reasoning and responses. Gupta et al. (2023) show how different personas, reflecting varied socio-demographic backgrounds, can induce biases in LLM responses. Similarly, Deshpande et al. (2023) demonstrate that persona variations can dramatically alter the toxicity levels in model outputs. This study also highlights the risks and challenges in deploying LLMs across diverse user interactions in persona-assigned LLMs.
CoT prompting for not only improving reasoning capabilities but also for mitigating biases in LLMs has been a recent development in the field. Kaneko et al. (2024) found that CoT prompting can effectively reduce gender bias in LLMs and that by adding step-by-step reasoning to model predictions, CoT prompting makes the prediction process clearer and cuts down on oft-repeated gender biases in LLMs.
3 Dataset
We use two bias detection datasets to measure the behavior of the LLMs, StereoSet (Nadeem et al., 2020) and Kamruzzaman et al.’s (Kamruzzaman et al., 2023) bias detection dataset. StereoSet covers gender, race, religion, and profession biases, specifically focusing on stereotypes related to these categories. Kamruzzaman et al.’s bias detection dataset covers age, beauty, institution, beauty in profession, and nationality bias and how LLMs associate specific attributes in these categories to unrelated general positive and negative descriptors. We will call Kamruzzaman et al.’s dataset the GenAssocBias dataset for the remainder of this paper.
Altogether, we evaluate LLMs over 9 types of social biases. We use the intrasentence subset of StereoSet as it aligns with the structure of GenAssocBias, leading to a uniform experimental design and analysis. The ‘profession’ bias category from StereoSet represents general stereotypical biases related to professions (e.g., The chess player was Asian) whereas the ‘beauty in profession’ from GenAssocBias represents the relationship between beauty and the status of a profession (e.g., She looks unattractive; she must be a street vendor).
4 Experimental Setup
We evaluate five major language models in this paper: 1) GPT-4 (Achiam et al., 2023), using the GPT-4 checkpoint on the OpenAI API; [^2] Due to the cost of running GPT-4, we use a sample of 2,100 entries from the GenAssocBias dataset to run GPT-4. We ensured that this sample contains an equal number of examples from each bias category. 2) GPT-3.5 (Brown et al., 2020), using the GPT-3.5-TURBO-INSTRUCT checkpoint on the OpenAI API; 3) Llama-2-7B (Touvron et al., 2023), using the TheBloke/Llama-2-7B-chat-GGUF [^3] We use 4-bit quantized Llama 2 and Mistral models due to availability of local computing resources. checkpoint on Huggingface; 4) Mistral-7B (Jiang et al., 2023), using the TheBloke/Mistral-7B-Instruct-v0.1-GGUF checkpoint on Huggingface; 5) Gemini (Team et al., 2023), using gemini-1.0-pro-latest on the Google Gemini API. Detailed procedures for the experiment are included in our supplementary materials, which also contain scripts for replication.
We use 12 different types of prompting techniques in our paper including the combinations of CoT, System 1, System 2, and Persona.
Type Prompt
Standard Prompt Fill in the BLANK with the most suitable word from the provided options…
CoT Let’s think step-by-step and fill in the BLANK. Given the context below, choose the most suitable word from the provided options …
System 1 You answer questions quickly. Given the context below, choose the most suitable word from the provided options …
System 2 You answer questions slowly and thoughtfully. Your answers are effortful and reliable. Given the context below, …
HP + System 1 Adopt the identity of a person who answers questions quickly. Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …
HP + System 2 Adopt the identity of a person who answers questions slowly and thoughtfully. Their answers are effortful and reliable. Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …
HP + CoT Adopt the identity of a person who answers questions by thinking them “step-by-step". Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …
HP + System 2 + CoT Adopt the identity of a person who answers questions slowly and thoughtfully. Their answers are effortful and reliable. They answer questions by thinking about them “step-by-step". Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …
MP + System 1 Adopt the identity of a machine that answers questions quickly. Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …
MP + System 2 Adopt the identity of a machine that answers questions slowly and thoughtfully. Its answers are effortful and reliable. Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …
MP + CoT Adopt the identity of a machine that answers questions by analyzing them “step-by-step". Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …
MP + System 2 + CoT Adopt the identity of a machine that answers questions slowly and thoughtfully. Its answers are effortful and reliable. It answers questions by analyzing them “step-by-step". Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …
Table 1: The types of prompts we use in our experiments. The “…” after each prompt are the common instructions which include “select the appropriate word from the options provided, without considering the order in which the options come in.” We provide options randomly to eliminate ordering bias and add “don’t add option number as well in your response” to minimize response postprocessing complexity.
4.1 Baseline Standard Prompt
We use the standard zero-shot prompting technique as our baseline. In this prompting, we ask questions directly without any special instructions. For example, we use the prompts “Fill in the BLANK with the most suitable word from the provided options…” This simple question allows the language model to choose a word that it thinks best fits the blank based on the given options.
4.2 CoT Prompt
For our CoT prompting, we use the zero-shot variant which includes the phrase “Let’s think step-by-step” in the prompt (Kaneko et al., 2024). This method encourages the LLM to unfold its thinking process in a structured and detailed manner. This is the simplest variant of the prompting method often compared to System 2 reasoning.
4.3 System 1 and System 2 Prompts
System 1 prompting encourages the model to make fast decisions, similar to the quick decision-making process in human cognition known as System 1. The System 1 prompt is “You answer questions quickly. Given the context below, choose the most suitable word from the provided options…”. This approach prompts the model to respond rapidly without deep deliberation, mimicking the intuitive human cognitive response. For System 2, which is characterized by slow, deliberate, and reliable decision-making, we use the prompt: “You answer questions slowly and thoughtfully. Your answers are effortful and reliable. Given the context below, ….”. This prompt encourages the model to process information deeply and consider multiple factors before making a decision. The aim is to simulate a more reflective thinking process in models. By promoting such processing, we intend to explore how the reduction of cognitive shortcuts—or the modeling of reduced cognitive shortcuts by the LLM—can decrease the reproduction of societal biases in LLM outputs.
4.4 Human and Machine Persona Prompts
In order to differentiate between the effects of dual process theory prompts on inherent LLM processing and the LLM’s model of human reasoning patterns, we incorporate prompting variants for human and machine personas. This is integrated with the other prompting methods (Standard, CoT, and Systems 1 and 2). So we add either a ‘Human Persona’ or a ‘Machine Persona’ by including the phrase ‘Adopt the identity of [persona]’, which influences how the LLM answers the following question. For instance, the ‘Human Persona with System 1’ (HP System 1) prompt is: ‘Adopt the identity of a person who answers questions quickly. Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …’. Similarly, the ‘Machine Persona with System 2’ (MP System 2) prompt is ‘Adopt the identity of a machine that answers questions slowly and thoughtfully. Its answers are effortful and reliable. Fill in the BLANK while staying in strict accordance with the nature of this identity. Given the context below, …’. See Table 1 for all the prompts we explore in this paper and how they realize persona, cognitive system, and CoT combinations. These varied personas help us explore how mimicking human-like cognitive processes in models might reduce inherent social biases.
5 Results & Analysis
Figure 2: Stereotypical Responses for each prompt, average across all the models and bias types. Here, MP stands for Machine Persona, HP stands for Human Persona.
Figure 3: Results with Standard Prompts and best-performing (in terms of least stereotypical engagement) prompts for each bias category and all the LLMs. Here, MP stands for Machine Persona, HP stands for Human Persona.
We present our main results in terms of stereotypical engagement/response rates, indicating the percentage (%) of responses that aligned with stereotypical judgments.
Overall Prompting Effects.
We present our overall stereotypical response rate for each prompt, averaged across all 5 models and 9 bias categories in Figure 2. Figure 2 shows that on average a Human Persona with System 2 prompting best reduces social bias in LLMs. We also see that on average the System 1 prompting is more stereotypical than other prompting techniques. This aligns with the behavior we would expect from dual process theory assuming that LLMs roughly model human cognition and that our System 1 and System 2 prompts appropriately trigger similar biases to the corresponding human cognitive systems.
A surprising result is that CoT prompting does not show any reduction in bias. In fact, for every available minimal pair in prompts we consider (Standard vs. CoT, HP System 2 vs. HP System 2 + CoT, and MP System 2 vs. MP System 2 + CoT), CoT leads to an increase in stereotypical responses.
Another surprising result is the effect of personas in prompts and how they relate to System 1 and System 2 prompts. First, we see that no matter which persona we use (Human or Machine) the stereotypical response rate drops (compare System 1 vs HP System 1 and MP System 1; make a similar comparison for System 2). At its face, this result is in direct contradiction with Gupta et al.’s (Gupta et al., 2023) and Deshpande et al. (2023) results where personas led to an increase in bias and toxicity. Our results reveal a more complicated situation. That is, adding a persona does not always lead to more bias in LLMs. Rather, LLM behavior changes based on the selected persona that the LLM is asked to model. In some cases, such as in the prior work, where the persona is a member of an underrepresented population, bias is revealed through its behavior. In our case, when the persona is a generic person or machine, it pushes the model away from socially biased responses. This suggests that having an LLM model a separate entity (human or machine) leads to more objective outputs, similar to Solomon’s Paradox observed in humans—this paradox is the observation that people reason more wisely regarding other’s problems than their own and that self-distancing may improve objective reasoning capabilities (Grossmann and Kross, 2014).
Things do not stop being interesting even yet. We find that when System 1 and System 2 prompts are combined with a human persona, their effects on social bias are amplified. That is the difference between the System 1 and System 2 responses is greater with the Human Persona + System 2 prompts having the least stereotypical responses overall. This combination results in a reduction of over 3% from the standard zero-shot prompt. While the Machine Persona leads to a reduction in bias, the difference in System 1 and System 2 results remains similar to the no-persona prompts. This suggests that while the LLM intrinsically models the two systems in dual process theory to some degree, its model of human cognition specifically has an even more exaggerated difference in these cognitive systems.
Model- and Bias-specific Prompting Effects.
All of the standard prompting results alongside the best performing prompting technique results for each bias category and model combination are presented in Figure 3. Here we see that the Human Persona with the System 2 (HP System 2) prompting technique often yields the least stereotypical responses, but that is not universal across models and bias categories. It outperforms all other prompting techniques in 13 cases. Similarly, the Human Persona in conjunction with System 2 and CoT (HP System 2 + CoT) prompting outperforms other prompting techniques in 9 cases. Only in one case, for profession bias and the GPT-4 model, the standard prompt outperform any other prompting techniques.
We also see from Figure 3 that the two open-weight models, Llama 2 and Mistral 7B, often have similar behaviors (with the exception of the religion category) and that the two more recent closed-weight models, GPT-4 and Gemini have similar behaviors. GPT-4 stands out in having the most cases where prompting variants make minimal improvement from the standard zero-shot prompt. This suggests that OpenAI has done some engineering on this front. Though it is unclear whether this is behind-the-scenes prompt modifications, analogous instruction finetuning of the model, or some other method to make the model robust to prompt variants. Next, we focus on each bias category.
Ageism.
Ageism bias in Figure 3 illustrates that various prompting techniques improved the performance of all models, reducing stereotypical responses by approximately 2 to 5 percent across the board compared to standard prompting.
Beauty.
From Figure 3, it is evident that using various prompting techniques improves the outcomes across all models, achieving up to 13 percent reduction in stereotypical responses for beauty bias. Notably, the Mistral 7B model exhibited the most promising results with the Human Persona System 2 (HP System 2) prompting technique. Furthermore, among the five models tested, four showed enhanced performance when prompted with HP System 2.
Beauty in Profession.
In the results depicted in Figure 3, it is evident that employing various prompting techniques leads to improved outcomes across all models, resulting in up to a 9 percent reduction in stereotypical responses for beauty in profession bias. Particularly noteworthy is the performance of the Gemini model when utilizing System 2 prompting, which shows the most promising results in reducing beauty in profession bias.
Gender.
We also observe a gender bias improvement of up to 9 percent across various models. For instance, the HP System 2 + CoT prompting technique performs well with GPT-3.5. Additionally, for GPT-4 and Gemini, the HP System 2 prompting yields the best performance. Notably, with Mistral 7B, HP System 1 outperforms other techniques. So, System 1 with Human Persona also reduces bias.
Institutional.
We observed a reduction in institutional bias across all models, although the percentage decrease was smaller compared to reductions in gender or beauty biases. Specifically, with Llama 2, we achieved around a 5 percent improvement when using HP System 1 for prompting.
Nationality.
Regarding nationality bias, the overall pattern of reduction is consistent across all models, similar to other biases. The combination of HP System 2 with CoT delivers the best performance for GPT-3.5 and Mistral 7B. For GPT-4 and Llama 2, the best results are achieved using HP alongside CoT.
Profession.
We achieved up to a 10 percent reduction in bias for the profession category. All models except GPT-4 exhibited less bias compared to the results from standard prompts. However, for GPT-4, standard prompting performed the best. In the case of Llama 2 and Mistral 7B, the HP System 2 prompts were the most effective, whereas for GPT-3.5 and Gemini, the System 2 prompts yielded the best results.
Race.
We observe a reduction in racial biases across all models, although the decrease is relatively small for GPT-4, akin to that observed with standard prompting techniques. In contrast, the Mistral 7B model using HP System 2 with CoT prompting shows a bias reduction of approximately 9 percent.
Religion.
We achieved a reduction in religious bias by up to 8 percent. Additionally, we observed reductions across all models, although the decrease in the Gemini model was relatively minor, nearly identical to that seen with standard prompting. For GPT-3.5, MP System 2 exhibited the best performance. For Llama 2, CoT prompting delivered the least stereotypical responses, marking the only scenario where CoT prompting outperformed other prompting techniques.
6 Which Cognitive System Does CoT Best Model?
Prompting Techniques $\tau$ $p$ $H_{0}$?
CoT Vs Standard 0.4755 0.0 Reject
CoT Vs System 1 0.4580 0.0 Reject
CoT Vs System 2 0.4339 0.0 Reject
HP CoT Vs HP System 1 0.4639 0.0 Reject
HP CoT Vs HP System 2 0.4423 0.0 Reject
MP CoT Vs MP System 1 0.4556 0.0 Reject
MP CoT Vs MP System 2 0.4372 0.0 Reject
Table 2: Kendall’s $\tau$ test results averaged across all bias types and models. We use a significance level of $\alpha<0.05$ to reject the null hypothesis.
Now we further investigate whether CoT prompting is most similar to System 2 reasoning in LLMs. While Figure 2 shows that the stereotyping rate of CoT is most similar to System 1 prompts, these may be from different test items. Here we tackle this question directly by computing the Kendall $\tau$ coefficient (Kendall, 1938) between CoT-prompted responses and those of the other variants. We selected the Kendall $\tau$ ranked correlation because there is a natural order to anti-stereotypical, neutral, and stereotypical categorical values in our datasets. Table 2 lists these results. From this, we find that CoT prompting is most similar to the Standard zero-shot prompt, followed by System 1 prompting. In fact, it is most dissimilar to System 2 prompting. This pattern holds for the Human Persona and Machine Persona variants, where CoT is least correlated with the System 2 prompt variant.
This raises questions regarding how to interpret the effect that CoT prompting has on LLM behavior. While prior work has found that CoT prompting leads to better multi-step mathematical and formal reasoning capabilities (Wei et al., 2022; Yu et al., 2023; Wang et al., 2023), which aligns with System 2 in dual process theory, we do not see the same effect in social biases in this study.
7 Invalid LLMs Responses
We excluded certain examples due to the language models providing invalid responses. These models did not consistently choose from the three options provided. The invalid responses sometimes included phrases from the context sentence but not from the options list. In other instances, the responses were completely unrelated to both the context sentence and the options list, means out-of-context responses. Additionally, a few responses were merely numerical, ranging from 1 to 3. Some responses indicate that certain stereotypes are present in a sentence and state that promoting stereotypes is inappropriate. When calculating the prevalence of stereotypical responses, we consider these responses, which demonstrate awareness of stereotypes, as anti-stereotype responses.
8 Conclusion
Our study has contributed to the understanding and reduction of social biases in LLMs through prompting techniques grounded in dual process theory. By harnessing the cognitive frameworks of System 1 and System 2, as well as the incorporation of human-like personas, our research not only clarifies the role of these cognitive processes in LLMs but also demonstrates practical methods for reducing biases. Our findings reveal that System 2 prompts, particularly when combined with a Human Persona, consistently reduce stereotypical judgments across various social bias categories. This indicates a profound potential for combining analytical thought processes and personalized prompting to enhance the ethical performance of LLMs. Furthermore, our use of different models and bias datasets has allowed us to explore the diverse applications of these techniques, ensuring our results are robust and applicable across different contexts. These findings underscore the potential of sophisticated prompting strategies in enhancing the ethical aspects of AI, pointing towards a future where LLMs can assist in creating more inclusive digital environments. Through continued exploration and refinement of these methods, we anticipate further advancements in the responsible deployment of AI technologies.

### v3

@@
 Dual process theory posits that human cognition arises via two systems.
-System 1, which is a quick, emotional, and intuitive process, which is subject to cognitive biases, and System 2, a slow, onerous, and deliberate process.
+System 1, which is a quick, emotional, and intuitive process, which is subject to cognitive biases, and System 2, is a slow, onerous, and deliberate process.
 NLP researchers often compare zero-shot prompting in LLMs to System 1 reasoning and chain-of-thought (CoT) prompting to System 2.
 In line with this interpretation, prior research has found that using CoT prompting in LLMs leads to reduced gender bias.
-We investigate the relationship between bias, CoT prompting, and dual process theory in LLMs directly.
-We compare zero-shot, CoT, and a variety of dual process theory-based prompting strategies on two bias datasets spanning nine different social bias categories.
-We also use human and machine personas to determine whether the effects of dual process theory in LLMs are based on modeling human cognition or inherent to the system.
-We find that a human persona, System 2, and CoT prompting all tend to reduce social biases in LLMs, though the best combination of features depends on the exact model and bias category—resulting in up to a 13 percent drop in stereotypical judgments by an LLM.111 Datasets and code are available at https://github.com/kamruzzaman15/Prompting-Techniques-to-Reduce-Social-Bias-in-LLMs
+We investigate the relationship between bias, CoT prompting, a debiasing prompt, and dual process theory in LLMs directly.
+We compare zero-shot CoT, debiasing, and a variety of dual process theory-based prompting strategies on two bias datasets spanning nine different social bias categories.
+We incorporate human and machine personas to determine whether the effects of dual process theory in LLMs exist independent of explicit persona models or are based on modeling human cognition.
+We find that a human persona, debiasing, System 2, and CoT prompting all tend to reduce social biases in LLMs, though the best combination of features depends on the exact model and bias category—resulting in up to a 19 percent drop in stereotypical judgments by an LLM.111 Datasets and code are available at https://github.com/kamruzzaman15/Reduce-Social-Bias-in-LLMs
 1 Introduction
@@
 In recent years, large language models (LLMs) like GPT-4 (Achiam et al., 2023), ChatGPT (Brown et al., 2020), Llama 2 (Touvron et al., 2023) have revolutionized many aspects of technology and society.
-These models display remarkable linguistic capabilities, crafting responses that not only mimic human language but also exhibit a depth of understanding previously unattainable in machine learning (Karanjai and Shi, 2024).
+These models display remarkable linguistic capabilities, crafting responses that not only mimic human language but also exhibit a depth of understanding previously unattainable in automation (Karanjai and Shi, 2024).
 A notable advancement in enhancing the reasoning capabilities of LLMs has been the introduction of CoT prompting (Wei et al., 2022).
 By simulating step-by-step reasoning, CoT prompting helps LLMs achieve higher levels of clarity and accuracy in complex tasks, significantly reducing errors inherent in simpler prompt designs.
-However, despite these advancements, LLMs continue to struggle with embedded social biases.
-These biases show up in different ways, including stereotyping and biased answers, which raises questions regarding the ethical use of LLMs in real-life applications.
-These biases are difficult to identify and even more challenging to eliminate wholesale due to the complex and opaque inner workings of AI models, the flexible and nuanced nature of human language, and the culturally dependent social rules that accompany language use.
+Despite these advancements, LLMs continue to struggle with embedded social biases, which raises questions regarding the ethical use of LLMs in real-life applications.
+These biases are difficult to identify and even more challenging to eliminate due to the complex and opaque inner workings of LLMs, the flexible and nuanced nature of human language, and the culturally dependent social rules that accompany language use.
 This task of mitigating social biases in LLMs is paramount to ensuring fairness and inclusivity in AI-driven communication and decisions.
-Applying dual process theory, a well-established psychological framework, to recent advancements in AI systems has illuminated possible pathways to enhancing the reliability and ethical footprint of LLMs by identifying where LLM behavior align with and diverge from human cognitive processes.
+Applying dual process theory, a well-established psychological framework, to recent AI advancements illuminates possible pathways to enhancing the reliability and ethical footprint of LLMs by identifying where LLM generations align with and diverge from human cognitive processes.
 In this paper, we use dual process theory-based prompting strategies, comparing their efficacy across multiple categories of social bias from two bias datasets.
-Our approach uniquely incorporates both human-like and machine-like personas to examine whether the effects of these cognitive theories in LLMs are merely analogical to human cognition or intrinsic to the AI systems themselves.
-In addition, we compare these results to the use of CoT prompting to test whether this prompting technique aligns with System 2 reasoning in LLMs as some have claimed in the past.
+Our approach incorporates human- and machine-like personas to examine whether the effects of these cognitive theories in LLMs are dependent on explicit co-modeling of human cognition or always implicitly modeled.
+We follow-up on this analysis by examining interactions with debiasing prompts designed specifically for social bias reduction.
+Finally, we compare these results to the use of CoT prompting to test whether this prompting technique aligns with System 2 reasoning in LLMs as some have claimed in the past.
 Figure 1 shows an example of how the human persona with System 2 prompting reduces stereotypical engagement over standard prompting.
-When we use standard zero-shot prompting in Figure 1 (a), we see that the Llama 2 model responds with a stereotypical answer.
-But when we use the prompting of the human persona with System 2 in Figure 1 (b), it instead responds with an anti-stereotypical answer.
+When we use standard zero-shot prompting in Figure 1 (a), we see that Llama 2 responds with a stereotypical answer.
+When we use human persona with System 2 prompting in Figure 1 (b), it instead responds with an anti-stereotypical answer.
 This paper’s contributions are the following.
-  • We explore the effects of 12 different prompting techniques including CoT, System 1, System 2, and Persona, across nine distinct social bias categories (ageism, beauty, beauty with profession, gender, institutional, nationality, profession, race, religion) in 5 cutting-edge LLMs.
-  • We find that incorporating System 2 and Human Persona to prompts both reduce social biases.
-The combination of the two lead to the largest reduction in social bias when averaged across models and bias categories.
-  • In contradiction to the assumptions made by researchers in the past (Hagendorff et al., 2023), we find that CoT prompting does not behave similarly to prompts that directly model System 2 for social biases.
+  • We explore the effects of 12 different prompting techniques including CoT, System 1, System 2, and Persona, across nine distinct social bias categories (ageism, beauty, beauty with profession, gender, institutional, nationality, profession, race, religion) in 5 LLMs.
+This is followed up with 6 prompting variations incorporating explicit debiasing.
+  • We find that incorporating a human persona is critical for controlling for biases in LLMs.
+While System 2 and explicit debiasing slightly reduce stereotypical responses on their own, combining them with a human persona lead to substantial improvements and the largest reductions in bias when averaged across models and bias categories.
+  • In line with Nighojkar’s (2024) recent results in the reasoning domain, we find that CoT prompting does not behave similarly to prompts that directly model System 2 for social biases.
 In fact, the rate of stereotypical responses is closest between CoT and System 1 prompts in all Persona variants (none, Human, and Machine).
+This contradicts often stated assumptions by researchers in the past (Hagendorff et al., 2023).
 2 Related Work
@@
 Hagendorff et al.
-(2023) look into reasoning biases that are similar to humans in LLMs.
-It shows that as these models get bigger and more complex, they start making intuitive mistakes, like those found in human System 1 thinking.
-This trend shifts with the introduction of ChatGPT models, which effectively avoid these reasoning traps by employing chain-of-thought processes reminiscent of human System 2 thinking, even when such explicit reasoning is inhibited.
+(2023) look into human-like reasoning biases in LMs.
+They show that as these models became bigger and more complex, they began making intuitive mistakes, like those found in human System 1 thinking.
+This trend shifted with the introduction of ChatGPT models, which effectively avoid these reasoning traps by employing chain-of-thought processes reminiscent of human System 2 thinking, even when such explicit reasoning is inhibited.
 In a significant enhancement to dual-process interaction within LLMs, Lin et al.
@@
 It shows its ability to efficiently solve complicated interactive challenges with less computational needs.
-Furthermore, Wei et al.
-(2022) and Sun (2024) contribute to our understanding of how embedding systematic reasoning in LLMs, through techniques like chain-of-thought prompting, can substantially improve performance across various cognitive tasks including complex arithmetic and symbolic reasoning.
-Coming to the Dual Process Theory, which is a central idea in understanding human thinking and decision-making.
-It distinguishes between two different kinds of thinking: fast, automatic (System 1), and slow, effortful (System 2).
+Coming to the debiasing studies, Furniturewala et al.
+(2024) investigate the use of structured prompting techniques for debiasing language models.
+The study explores three categories of prompts: Prefix Prompting, Self-Refinement, and Implication Prompting.
+Chisca et al.
+(2024) focus on mitigating biases in LLMs through prompt-tuning.
+This approach introduces small sets of trainable token embeddings that are concatenated to input sequences, aiming to reduce biases without major modifications to the model’s architecture.
+Dual Process Theory is a psychological account of how human thinking and decision-making arise from two distinct modes.
+It distinguishes between fast, automatic (System 1), and slow, effortful (System 2) modes of thinking.
 System 1 enables quick comprehension through associations and pre-existing knowledge.
@@
 These systems guide our reasoning, decision-making, and learning processes in various cognitive tasks (Frankish, 2010).
-The theory illuminates the intricate relation between intuitive, heuristic thinking and analytical, rule-based cognition (Evans and Stanovich, 2013).
-In language understanding, Dual Process Theory explains how we can quickly and easily understand the general meanings of language input, but also switch to a more deliberate process when we encounter more complex language use, whether that be, for example, complex embedded syntactic structures, ambiguity in meaning, or uncertain pragmatic contexts.
-This shows how both quick thinking and more careful thinking work together (Ferreira and Huettig, 2023).
+The theory illuminates the intricate relation between intuitive, heuristic thinking and analytical, rule-based cognition (Evans and Stanovich, 2013; Ferreira and Huettig, 2023).
 Our understanding of our own thinking and knowing our mind’s state is connected to this two-part idea.
@@
 They can also come from a fight between heuristic and logical intuitions that happen at the same time.
-What happens next depends on which intuition is stronger and if the person realizes there is a conflict.
-If the relevant logical intuition is available, a correct judgment can be made without relying on the more resource-intensive System 2 cognitive process.
 This shows that logical thinking does not just belong to System 2 (De Neys and Pennycook, 2019).
 These more recent developments in the theory reveal a more nuanced picture of the ideal cognitive system selection for any given task.
-Building on this theory, Bellini-Leite (2023) discusses the implementation of System 2 analytical processes in LLMs to mimic careful human reasoning, potentially mitigating frequent errors and enhancing decision-making capabilities in these models.
+Bellini-Leite (2023) discusses how methods such as CoT and tree-of-thought prompting in LLMs are suggestive of System 2 human reasoning, potentially mitigating frequent errors and enhancing reliability in these models.
+Nighojkar (2024) go on to test this correspondence by comparing against results from human experiments.
+He finds that CoT prompting does not simply mimic System 2.
+Rather, it leads to better agreement with human responses in both System 1 and System 2 triggering instructions.
 Recent research on LLMs has found that assigning personas to LLMs can notably impact their reasoning and responses.
-Gupta et al.
-(2023) show how different personas, reflecting varied socio-demographic backgrounds, can induce biases in LLM responses.
-Similarly, Deshpande et al.
-(2023) demonstrate that persona variations can dramatically alter the toxicity levels in model outputs.
-This study also highlights the risks and challenges in deploying LLMs across diverse user interactions in persona-assigned LLMs.
-CoT prompting for not only improving reasoning capabilities but also for mitigating biases in LLMs has been a recent development in the field.
-Kaneko et al.
-(2024) found that CoT prompting can effectively reduce gender bias in LLMs and that by adding step-by-step reasoning to model predictions, CoT prompting makes the prediction process clearer and cuts down on oft-repeated gender biases in LLMs.
+Beck et al.
+(2024) highlights that sociodemographic prompting can significantly influence model predictions and improve zero-shot learning performance in subjective tasks.
+But the effectiveness of this approach varies across different models, dataset types, and sociodemographic factors.
+de Araujo and Roth (2024) explore how different personas, such as roles or identities (e.g., gender, occupation, or cultural background) influence the behavior and outputs of LLMs across various tasks.
+The research involves giving 162 unique personas to various LLMs and assessing how well they perform on different tasks and it turns out that personas not only affect task performance, bias, and model refusals but also show varied responses depending on the assigned persona.
+In addition to personas, explicitly debiasing instructions in prompts have been found to effectively reduce gender bias in LLMs (Kaneko et al., 2024).
 3 Dataset
-We use two bias detection datasets to measure the behavior of the LLMs, StereoSet (Nadeem et al., 2020) and Kamruzzaman et al.’s (Kamruzzaman et al., 2023) bias detection dataset.
+We use two bias detection datasets to measure the behavior of the LLMs, StereoSet (Nadeem et al., 2020) and Kamruzzaman et al.’s (Kamruzzaman et al., 2024) bias detection dataset.
 StereoSet covers gender, race, religion, and profession biases, specifically focusing on stereotypes related to these categories.
-Kamruzzaman et al.’s bias detection dataset covers age, beauty, institution, beauty in profession, and nationality bias and how LLMs associate specific attributes in these categories to unrelated general positive and negative descriptors.
-We will call Kamruzzaman et al.’s dataset the GenAssocBias dataset for the remainder of this paper.
+Kamruzzaman et al.’s bias detection dataset covers age, beauty, institution, beauty in the profession, and nationality bias and how LLMs associate specific attributes in these categories to unrelated general positive and negative descriptors.
+We will call this dataset the GenAssocBias dataset for the remainder of this paper.
 Altogether, we evaluate LLMs over 9 types of social biases.
@@
 We ensured that this sample contains an equal number of examples from each bias category.
-2) GPT-3.5 (Brown et al., 2020), using the GPT-3.5-TURBO-INSTRUCT checkpoint on the OpenAI API; 3) Llama-2-7B (Touvron et al., 2023), using the TheBloke/Llama-2-7B-chat-GGUF [^3] We use 4-bit quantized Llama 2 and Mistral models due to availability of local computing resources.
+2) GPT-3.5 (Brown et al., 2020), using the GPT-3.5-TURBO-INSTRUCT checkpoint on the OpenAI API; 3) Llama-2-13B (Touvron et al., 2023), using the TheBloke/Llama-2-13B-chat-GGUF [^3] We use 4-bit quantized Llama 2 and Mistral models due to limited local computing resources.
 checkpoint on Huggingface; 4) Mistral-7B (Jiang et al., 2023), using the TheBloke/Mistral-7B-Instruct-v0.1-GGUF checkpoint on Huggingface; 5) Gemini (Team et al., 2023), using gemini-1.0-pro-latest on the Google Gemini API.
-Detailed procedures for the experiment are included in our supplementary materials, which also contain scripts for replication.
-We use 12 different types of prompting techniques in our paper including the combinations of CoT, System 1, System 2, and Persona.
+We use 12 different types of prompting techniques in the main experiments including the combinations of CoT, System 1, System 2, and Persona.
+This is followed up with 6 additional prompting techniques with explicit debiasing.
 Type Prompt
@@
 Given the context below, …
-Table 1: The types of prompts we use in our experiments.
+Table 1: The types of prompts we use in our main experiments.
 The “…” after each prompt are the common instructions which include “select the appropriate word from the options provided, without considering the order in which the options come in.” We provide options randomly to eliminate ordering bias and add “don’t add option number as well in your response” to minimize response postprocessing complexity.
+Type Prompt
+Standard Prompt + Debias Fill in the BLANK with the most suitable word from the provided options. The response should be unbiased and does not rely on stereotypes…
+HP + Debias Adopt the identity of a person who always makes sure that their answer is unbiased and does not rely on stereotypes. Fill in the BLANK …
+System 2 + Debias You answer questions slowly and thoughtfully.
+Your answers are effortful and reliable.
+You always make sure that your answer is unbiased and does not rely on stereotypes.
+Fill in the BLANK …
+HP + System 2 + Debias Adopt the identity of a person who answers questions slowly and thoughtfully.
+Their answers are effortful and reliable.
+They always make sure that their answer is unbiased and does not rely on stereotypes.
+Fill in the …
+CoT + Debias Let’s think step-by-step and fill in the BLANK. Make sure that the answer is unbiased and does not rely on stereotypes …
+HP + System 2 + CoT + Debias Adopt the identity of a person who answers questions slowly and thoughtfully.
+Their answers are effortful and reliable.
+They always make sure that their answer is unbiased and does not rely on stereotypes.
+They answer questions by thinking about them “step-by-step".
+Fill …
+Table 2: The types of prompts we use when incorporating debiasing explicitly. The “…” represent the same as in Table 1
 4.1 Baseline Standard Prompt
@@
 4.3 System 1 and System 2 Prompts
-System 1 prompting encourages the model to make fast decisions, similar to the quick decision-making process in human cognition known as System 1.
+System 1 prompting frames the context so that LLMs generate text reflecting the quick decision-making process in human cognition known as System 1.
 The System 1 prompt is “You answer questions quickly.
 Given the context below, choose the most suitable word from the provided options…”.
-This approach prompts the model to respond rapidly without deep deliberation, mimicking the intuitive human cognitive response.
-For System 2, which is characterized by slow, deliberate, and reliable decision-making, we use the prompt: “You answer questions slowly and thoughtfully.
+This approach prompts the model to generate text mimicking the intuitive human cognitive response which does not include deep deliberation.
+For modeling System 2, characterized by slow, deliberate, and reliable decision-making, we use the prompt: “You answer questions slowly and thoughtfully.
 Your answers are effortful and reliable.
 Given the context below, ….”.
-This prompt encourages the model to process information deeply and consider multiple factors before making a decision.
-The aim is to simulate a more reflective thinking process in models.
-By promoting such processing, we intend to explore how the reduction of cognitive shortcuts—or the modeling of reduced cognitive shortcuts by the LLM—can decrease the reproduction of societal biases in LLM outputs.
+This prompt frames the context to generate text that reflects human responses processing information deeply and considering multiple factors.
+By using this context, we explore how the approximation of reduced cognitive shortcuts by the LLM can decrease the reproduction of societal biases in LLM outputs.
 4.4 Human and Machine Persona Prompts
-In order to differentiate between the effects of dual process theory prompts on inherent LLM processing and the LLM’s model of human reasoning patterns, we incorporate prompting variants for human and machine personas.
+In order to differentiate between the effects of dual process theory prompts on bare LLM processing and the LLM’s model of human reasoning patterns, we incorporate prompting variants for human and machine personas.
 This is integrated with the other prompting methods (Standard, CoT, and Systems 1 and 2).
-So we add either a ‘Human Persona’ or a ‘Machine Persona’ by including the phrase ‘Adopt the identity of [persona]’, which influences how the LLM answers the following question.
+We add either a ‘Human Persona’ or a ‘Machine Persona’ by including the phrase ‘Adopt the identity of [persona]’, which influences how the LLM answers the following question.
 For instance, the ‘Human Persona with System 1’ (HP System 1) prompt is: ‘Adopt the identity of a person who answers questions quickly.
@@
 Figure 2: Stereotypical Responses for each prompt, average across all the models and bias types. Here, MP stands for Machine Persona, HP stands for Human Persona.
+Figure 3: Stereotypical Responses for the debiasing prompt follow-up experiment (orange colored). The blue colored bars are anchors from Figure 2 for easy comparison.
 Figure 4: Results with Standard Prompts and best-performing (in terms of least stereotypical engagement) prompts for each bias category and all the LLMs.
@@
 MP System 2 + CoT), CoT leads to an increase in stereotypical responses.
-Another surprising result is the effect of personas in prompts and how they relate to System 1 and System 2 prompts.
+Another result is the effect of personas in prompts and how they relate to System 1 and System 2 prompts.
 First, we see that no matter which persona we use (Human or Machine) the stereotypical response rate drops (compare System 1 vs HP System 1 and MP System 1; make a similar comparison for System 2).
-At its face, this result is in direct contradiction with Gupta et al.’s (Gupta et al., 2023) and Deshpande et al.
-(2023) results where personas led to an increase in bias and toxicity.
-Our results reveal a more complicated situation.
-That is, adding a persona does not always lead to more bias in LLMs.
-Rather, LLM behavior changes based on the selected persona that the LLM is asked to model.
-In some cases, such as in the prior work, where the persona is a member of an underrepresented population, bias is revealed through its behavior.
-In our case, when the persona is a generic person or machine, it pushes the model away from socially biased responses.
-This suggests that having an LLM model a separate entity (human or machine) leads to more objective outputs, similar to Solomon’s Paradox observed in humans—this paradox is the observation that people reason more wisely regarding other’s problems than their own and that self-distancing may improve objective reasoning capabilities (Grossmann and Kross, 2014).
-Things do not stop being interesting even yet.
-We find that when System 1 and System 2 prompts are combined with a human persona, their effects on social bias are amplified.
-That is the difference between the System 1 and System 2 responses is greater with the Human Persona + System 2 prompts having the least stereotypical responses overall.
+This suggests that having an LLM model as a separate entity (human or machine) leads to less socially biased outputs.
+When System 1 and System 2 prompts are combined with a human persona, their effects on social bias are amplified.
+The difference between the System 1 and System 2 responses is greater with the Human Persona + System 2 prompts having the least stereotypical responses overall.
 This combination results in a reduction of over 3% from the standard zero-shot prompt.
 While the Machine Persona leads to a reduction in bias, the difference in System 1 and System 2 results remains similar to the no-persona prompts.
-This suggests that while the LLM intrinsically models the two systems in dual process theory to some degree, its model of human cognition specifically has an even more exaggerated difference in these cognitive systems.
-Model- and Bias-specific Prompting Effects.
+This suggests that while the LLM independent of a persona differentiates the two systems in dual process theory to some degree, its model of human cognition has an even more exaggerated difference in these cognitive systems.
+5.1 Debiasing Prompt Follow-up
+From Figure 2, we see that HP System 2 and HP System 2 with CoT prompting techniques perform substantially better than other prompt settings on average.
+We perform a follow-up experiment based on these two techniques, investigating explicitly debiasing prompts, similar to Kaneko et al.
+(2024).
+We add 6 debiasing prompting techniques: variations of HP, System 2, and CoT prompt combinations.
+The exact debasing prompts are shown in Table 2.
+Figure 3 shows the overall stereotypical response rates for these debiasing-incorporated prompting techniques averaged across all five models and 9 bias categories.
+It shows that the HP Debias prompt performs best compared to all other techniques.
+Similar to System 2 prompts, we find that the bias reduction effects of explicit debiasing is amplified by a human persona.
+While System 2 and CoT are complementary with explicit debiasing, they do not provide additional benefit beyond the HP and debiasing combination.
+In fact, we find that on average other prompt features interfere with and degrade the best performing HP Debias setting.
+This suggests that HP System 2 and HP Debias prompting achieve similar changes in model generations but with Debias having a more social focus.
+This is not surprising when we look at the prompts.
+The debias prompt can be seen as a brief social-oriented structured reasoning prompt.
+5.2 Model- and Bias-specific Prompting Effects
+We now turn to specific model-bias category combinations.
 All of the standard prompting results alongside the best performing prompting technique results for each bias category and model combination are presented in Figure 4.
-Here we see that the Human Persona with the System 2 (HP System 2) prompting technique often yields the least stereotypical responses, but that is not universal across models and bias categories.
-It outperforms all other prompting techniques in 13 cases.
-Similarly, the Human Persona in conjunction with System 2 and CoT (HP System 2 + CoT) prompting outperforms other prompting techniques in 9 cases.
-Only in one case, for profession bias and the GPT-4 model, the standard prompt outperform any other prompting techniques.
-We also see from Figure 3 that the two open-weight models, Llama 2 and Mistral 7B, often have similar behaviors (with the exception of the religion category) and that the two more recent closed-weight models, GPT-4 and Gemini have similar behaviors.
+Here we see that the Human Persona with the System 2 (HP System 2) and Human Persona with debias (HP Debias) prompting technique often yields the least stereotypical responses, but that is not universal across models and bias categories.
+HP Debias outperforms all other prompting techniques in 14 cases.
+Similarly, the Human Persona in conjunction with System 2 (HP System 2) prompting outperforms other prompting techniques in 11 cases.
+Only in one case, for profession bias and the GPT-4 model, the standard prompt outperforms any other prompting techniques.
+We also see from Figure 4 that the two open-weight models, Llama 2 and Mistral 7B, often have similar behaviors and that the two more recent closed-weight models, GPT-4 and Gemini-1.0 have similar behaviors.
 GPT-4 stands out in having the most cases where prompting variants make minimal improvement from the standard zero-shot prompt.
@@
 Ageism.
-Ageism bias in Figure 3 illustrates that various prompting techniques improved the performance of all models, reducing stereotypical responses by approximately 2 to 5 percent across the board compared to standard prompting.
+We see no consistent prompt setting that performs best on ageism. Stereotypical responses in models are reduced by 2 to 6 percent in the best prompt settings.
 Beauty.
-From Figure 3, it is evident that using various prompting techniques improves the outcomes across all models, achieving up to 13 percent reduction in stereotypical responses for beauty bias.
-Notably, the Mistral 7B model exhibited the most promising results with the Human Persona System 2 (HP System 2) prompting technique.
-Furthermore, among the five models tested, four showed enhanced performance when prompted with HP System 2.
+Prompt variants in our experiments show substantial improvements on beauty bias across all considered models—up to 19 percent reduction in stereotypical responses in Mistral-7B using the HP System 2 prompt.
+The remaining 4 models also show major improvements in beauty bias, all using the HP Debias prompt.
 Beauty in Profession.
-In the results depicted in Figure 3, it is evident that employing various prompting techniques leads to improved outcomes across all models, resulting in up to a 9 percent reduction in stereotypical responses for beauty in profession bias.
-Particularly noteworthy is the performance of the Gemini model when utilizing System 2 prompting, which shows the most promising results in reducing beauty in profession bias.
+Gemini-1.0 shows a 13 percent reduction in stereotypical responses for beauty in profession bias using System 2 Debias prompting.
+The best prompt setting is inconsistent for this bias category, with HP System 2 showing the largest bias reduction for GPT-4 and Llama 2 while HP Debias and HP System 2 Debias results in the largest bias reduction in GPT-3.5 and Mistral-7B, respectively.
 Gender.
-We also observe a gender bias improvement of up to 9 percent across various models.
-For instance, the HP System 2 + CoT prompting technique performs well with GPT-3.5.
-Additionally, for GPT-4 and Gemini, the HP System 2 prompting yields the best performance.
-Notably, with Mistral 7B, HP System 1 outperforms other techniques.
-So, System 1 with Human Persona also reduces bias.
+We see no consistent prompt setting that best reduces gender bias, but the best setting leads to consistent bias reductions.
+Interestingly, the open-weight models, Llama-2 and Mistral-7B show less stereotypical gender responses before explored prompt-based improvements than the other models after.
 Institutional.
-We observed a reduction in institutional bias across all models, although the percentage decrease was smaller compared to reductions in gender or beauty biases.
-Specifically, with Llama 2, we achieved around a 5 percent improvement when using HP System 1 for prompting.
+Again, we observe no consistent prompt setting that best reduces institutional bias.
+However, the percentage decrease was smaller compared to reductions in gender or beauty biases.
+With Llama 2, we achieved about a 5 percent improvement when using HP System 1 for prompting.
 Nationality.
-Regarding nationality bias, the overall pattern of reduction is consistent across all models, similar to other biases.
+Regarding nationality bias, the overall pattern of reduction is consistent across all models, similar to other biases, but the best prompting method differs.
 The combination of HP System 2 with CoT delivers the best performance for GPT-3.5 and Mistral 7B.
-For GPT-4 and Llama 2, the best results are achieved using HP alongside CoT.
+GPT-4 shows the least overall nationality bias, achieved using HP alongside CoT.
 Profession.
-We achieved up to a 10 percent reduction in bias for the profession category.
-All models except GPT-4 exhibited less bias compared to the results from standard prompts.
-However, for GPT-4, standard prompting performed the best.
-In the case of Llama 2 and Mistral 7B, the HP System 2 prompts were the most effective, whereas for GPT-3.5 and Gemini, the System 2 prompts yielded the best results.
+We achieved up to an 11 percent reduction in bias for the profession category.
+Here we see the single bias-model combination where standard prompting performed best in GPT-4.
+The other models all show substantial improvements.
+GPT-3.5 and Gemini-1.0, System 2 prompts yielded the best results.
+For Mistral-7B, Hp System 2 and for Llama 2, HP Debias were the best.
 Race.
@@
 In contrast, the Mistral 7B model using HP System 2 with CoT prompting shows a bias reduction of approximately 9 percent.
+The best-performing prompting technique for race is HP Debias for GPT-3.5 which reduces around 17 percent of stereotypical engagement.
+Here HP Debias was the best prompting technique in three of the five models.
 Religion.
-We achieved a reduction in religious bias by up to 8 percent.
-Additionally, we observed reductions across all models, although the decrease in the Gemini model was relatively minor, nearly identical to that seen with standard prompting.
-For GPT-3.5, MP System 2 exhibited the best performance.
-For Llama 2, CoT prompting delivered the least stereotypical responses, marking the only scenario where CoT prompting outperformed other prompting techniques.
-6 Which Cognitive System Does CoT Best Model?
+We achieved a reduction in religious bias by up to 13 percent.
+Additionally, we observed reductions across all models, although the decrease in the GPT-4 model was relatively minor.
+Again, we observe no consistent prompt setting that best reduces religious bias.
+6 Does CoT Prompting Best Model System 2?
 Prompting Techniques $\tau$ $p$ $H_{0}$?
-CoT Vs Standard 0.4755 0.0 Reject
-CoT Vs System 1 0.4580 0.0 Reject
-CoT Vs System 2 0.4339 0.0 Reject
-HP CoT Vs HP System 1 0.4639 0.0 Reject
-HP CoT Vs HP System 2 0.4423 0.0 Reject
-MP CoT Vs MP System 1 0.4556 0.0 Reject
-MP CoT Vs MP System 2 0.4372 0.0 Reject
+CoT Vs Standard 0.476 0.0 Reject
+CoT Vs System 1 0.458 0.0 Reject
+CoT Vs System 2 0.434 0.0 Reject
+HP CoT Vs HP System 1 0.464 0.0 Reject
+HP CoT Vs HP System 2 0.442 0.0 Reject
+MP CoT Vs MP System 1 0.456 0.0 Reject
+MP CoT Vs MP System 2 0.437 0.0 Reject
 Table 3: Kendall’s $\tau$ test results averaged across all bias types and models. We use a significance level of $\alpha<0.05$ to reject the null hypothesis.
-Now we further investigate whether CoT prompting is most similar to System 2 reasoning in LLMs.
+Now we further investigate whether CoT prompting is most similar to the way that LLMs model System 2 reasoning.
 While Figure 2 shows that the stereotyping rate of CoT is most similar to System 1 prompts, these may be from different test items.
 Here we tackle this question directly by computing the Kendall $\tau$ coefficient (Kendall, 1938) between CoT-prompted responses and those of the other variants.
-We selected the Kendall $\tau$ ranked correlation because there is a natural order to anti-stereotypical, neutral, and stereotypical categorical values in our datasets.
+We use the Kendall $\tau$ ranked correlation because there is a natural order to anti-stereotypical, neutral, and stereotypical categorical values in our datasets.
 Table 3 lists these results.
@@
 This pattern holds for the Human Persona and Machine Persona variants, where CoT is least correlated with the System 2 prompt variant.
-This raises questions regarding how to interpret the effect that CoT prompting has on LLM behavior.
-While prior work has found that CoT prompting leads to better multi-step mathematical and formal reasoning capabilities (Wei et al., 2022; Yu et al., 2023; Wang et al., 2023), which aligns with System 2 in dual process theory, we do not see the same effect in social biases in this study.
+Our study aligns with Nighojkar’s (2024) results showing that CoT does not specifically resemble System 2.
+Nighojkar (2024) found that CoT prompting leads to LLMs better modeling human behavior, whether that is System 1 or System 2 depending on which cognitive process the setting triggers.
+While prior work has found that CoT prompting leads to better multi-step mathematical and formal reasoning capabilities (Wei et al., 2022; Yu et al., 2023; Wang et al., 2023), that align with System 2 cognitive processes, the growing body of evidence suggests that this is because the formal reasoning setting contextualizes LLMs to generate text reflecting System 2 reasoning in people.
 7 Invalid LLMs Responses
@@
 Our study has contributed to the understanding and reduction of social biases in LLMs through prompting techniques grounded in dual process theory.
-By harnessing the cognitive frameworks of System 1 and System 2, as well as the incorporation of human-like personas, our research not only clarifies the role of these cognitive processes in LLMs but also demonstrates practical methods for reducing biases.
+By harnessing the cognitive frameworks of System 1 and System 2, as well as the incorporation of human-like personas and debiasing prompts, our research not only clarifies the role of these cognitive processes in LLMs but also demonstrates practical methods for reducing biases.
 Our findings reveal that System 2 prompts, particularly when combined with a Human Persona, consistently reduce stereotypical judgments across various social bias categories.
+Biases were further reduced when using a debiasing prompt, which can be seen as a social bias-focused System 2 prompt, along with a human persona.
 This indicates a profound potential for combining analytical thought processes and personalized prompting to enhance the ethical performance of LLMs.

### v4

@@
 System 1, which is a quick, emotional, and intuitive process, which is subject to cognitive biases, and System 2, is a slow, onerous, and deliberate process.
-NLP researchers often compare zero-shot prompting in LLMs to System 1 reasoning and chain-of-thought (CoT) prompting to System 2.
-In line with this interpretation, prior research has found that using CoT prompting in LLMs leads to reduced gender bias.
-We investigate the relationship between bias, CoT prompting, a debiasing prompt, and dual process theory in LLMs directly.
-We compare zero-shot CoT, debiasing, and a variety of dual process theory-based prompting strategies on two bias datasets spanning nine different social bias categories.
-We incorporate human and machine personas to determine whether the effects of dual process theory in LLMs exist independent of explicit persona models or are based on modeling human cognition.
-We find that a human persona, debiasing, System 2, and CoT prompting all tend to reduce social biases in LLMs, though the best combination of features depends on the exact model and bias category—resulting in up to a 19 percent drop in stereotypical judgments by an LLM.111 Datasets and code are available at https://github.com/kamruzzaman15/Reduce-Social-Bias-in-LLMs
+Prior research in LLMs found that using chain-of-thought (CoT) prompting in LLMs, which has been often compared to System 2 reasoning, can lead to reduced gender bias.
+Along these lines, we investigate the relationship between bias, CoT prompting, a direct debiasing, and dual process theory modeling in LLMs.
+We compare zero-shot CoT, debiasing, and dual process theory-based prompting strategies on two bias datasets spanning nine different social bias categories.
+We incorporate human and machine personas to determine whether LLM modeling of the effects of dual process theory exist independent of explicit persona models or are tied to the LLM’s modeling of human-like generation.
+We find that a human persona, debiasing, System 2, and CoT prompting all tend to reduce social biases in LLMs, though the best combination of features depends on the exact model and bias category—resulting in up to a 33 percent drop in stereotypical judgments by an LLM.
+[^1] Our code is available at https://github.com/kamruzzaman15/Reduce-Social-Bias-in-LLMs.
+00footnotetext: This work has been accepted at RANLP-2025.
 1 Introduction
-Figure 1: Example of Standard Prompting and Human Persona with System 2 Prompting for Llama 2 model in the race bias category
-In recent years, large language models (LLMs) like GPT-4 (Achiam et al., 2023), ChatGPT (Brown et al., 2020), Llama 2 (Touvron et al., 2023) have revolutionized many aspects of technology and society.
+Figure 1: Example of Standard Prompting and Human Persona with System 2 Prompting for Llama3.3 model in the race bias category
+In recent years, large language models (LLMs) like GPT-4 (Achiam et al., 2023), ChatGPT (Brown et al., 2020), Llama (Touvron et al., 2023) have revolutionized many aspects of technology and society.
 These models display remarkable linguistic capabilities, crafting responses that not only mimic human language but also exhibit a depth of understanding previously unattainable in automation (Karanjai and Shi, 2024).
@@
 This task of mitigating social biases in LLMs is paramount to ensuring fairness and inclusivity in AI-driven communication and decisions.
+Previous approaches to bias mitigation in LLMs often rely on fine-tuning techniques, which require access to the model’s weights and training mechanisms Zmigrod et al.
+(2019); Liang et al.
+(2021); Schick et al.
+(2021).
+While effective, these methods are computationally expensive and impractical for many state-of-the-art LLMs that remain closed-source or available only through restricted APIs.
+As an alternative, prompting strategies offer a lightweight and accessible method for steering model outputs toward fairness.
+By leveraging well-motivated and experimentally validated prompts, end-users can reduce bias in a manner that is practical for resource-constrained scenarios and effective for closed models.
 Applying dual process theory, a well-established psychological framework, to recent AI advancements illuminates possible pathways to enhancing the reliability and ethical footprint of LLMs by identifying where LLM generations align with and diverge from human cognitive processes.
 In this paper, we use dual process theory-based prompting strategies, comparing their efficacy across multiple categories of social bias from two bias datasets.
-Our approach incorporates human- and machine-like personas to examine whether the effects of these cognitive theories in LLMs are dependent on explicit co-modeling of human cognition or always implicitly modeled.
+Our approach incorporates human- and machine-like personas to examine whether the effects of these cognitive theories in the generations of LLMs are dependent on explicit co-modeling of human cognitive patterns or always implicitly modeled.
 We follow-up on this analysis by examining interactions with debiasing prompts designed specifically for social bias reduction.
-Finally, we compare these results to the use of CoT prompting to test whether this prompting technique aligns with System 2 reasoning in LLMs as some have claimed in the past.
 Figure 1 shows an example of how the human persona with System 2 prompting reduces stereotypical engagement over standard prompting.
-When we use standard zero-shot prompting in Figure 1 (a), we see that Llama 2 responds with a stereotypical answer.
-When we use human persona with System 2 prompting in Figure 1 (b), it instead responds with an anti-stereotypical answer.
 This paper’s contributions are the following.
+  • To the best of our knowledge, we are the first to investigate the intersection of dual process theory, and social bias in LLMs.
+We incorporate System 1 and System 2 with persona to reduce social biases in LLMs.
   • We explore the effects of 12 different prompting techniques including CoT, System 1, System 2, and Persona, across nine distinct social bias categories (ageism, beauty, beauty with profession, gender, institutional, nationality, profession, race, religion) in 5 LLMs.
@@
   • We find that incorporating a human persona is critical for controlling for biases in LLMs.
-While System 2 and explicit debiasing slightly reduce stereotypical responses on their own, combining them with a human persona lead to substantial improvements and the largest reductions in bias when averaged across models and bias categories.
-  • In line with Nighojkar’s (2024) recent results in the reasoning domain, we find that CoT prompting does not behave similarly to prompts that directly model System 2 for social biases.
-In fact, the rate of stereotypical responses is closest between CoT and System 1 prompts in all Persona variants (none, Human, and Machine).
-This contradicts often stated assumptions by researchers in the past (Hagendorff et al., 2023).
+While System 2 prompting and explicit debiasing slightly reduce stereotypical responses on their own, combining them with a human persona lead to substantial improvements and the largest reductions in bias when averaged across models and bias categories.
 2 Related Work
-Recent studies have explored how reasoning in LLMs can exhibit biases similar to human cognitive processes.
+Human-Like Reasoning Biases in LLMs.
+Recent studies have explored how reasoning in LLMs can exhibit biases similar to human cognitive processes Suri et al.
+(2024); Macmillan-Scott and Musolesi (2024); Ando et al.
+(2023).
 Hagendorff et al.
-(2023) look into human-like reasoning biases in LMs.
-They show that as these models became bigger and more complex, they began making intuitive mistakes, like those found in human System 1 thinking.
-This trend shifted with the introduction of ChatGPT models, which effectively avoid these reasoning traps by employing chain-of-thought processes reminiscent of human System 2 thinking, even when such explicit reasoning is inhibited.
-In a significant enhancement to dual-process interaction within LLMs, Lin et al.
-(2024) introduce SWIFTSAGE a new dual-module framework for better action planning in complex interactive reasoning tasks.
-This framework combines behavior cloning and prompting large language models.
-It includes the SWIFT module for quick, intuitive responses, and the SAGE module for careful, detailed planning.
-Tested on 30 tasks in the ScienceWorld benchmark, SWIFTSAGE greatly outperforms current methods like SayCan, ReAct, and Reflexion.
-It shows its ability to efficiently solve complicated interactive challenges with less computational needs.
-Coming to the debiasing studies, Furniturewala et al.
-(2024) investigate the use of structured prompting techniques for debiasing language models.
-The study explores three categories of prompts: Prefix Prompting, Self-Refinement, and Implication Prompting.
-Chisca et al.
-(2024) focus on mitigating biases in LLMs through prompt-tuning.
-This approach introduces small sets of trainable token embeddings that are concatenated to input sequences, aiming to reduce biases without major modifications to the model’s architecture.
+(2023) look into human-like reasoning biases in LLMs and find that as these models became bigger and more complex, they began making intuitive mistakes, like those found in human System 1 thinking.
+Moreover, studies have found that LLMs can replicate human-like cognitive biases, such as the representativeness heuristic, leading to stereotypical reasoning patterns Wang et al.
+(2024); Ryu et al.
+(2024).
+Debiasing Approaches in LLMs.
+Recent work on LLM debiasing explores both explicit and implicit strategies, including prompt-tuning, which embeds trainable tokens into input sequences to reduce bias without altering model architecture (Chisca et al., 2024).
+Another effective strategy involves self-diagnosis and self-debiasing, allowing models to recognize and reduce their own biases through controlled decoding mechanisms (Schick et al., 2021; Gallegos et al., 2024).
+Further refinements to these methods integrate assisted self-debiasing with external fairness constraints to guide LLM outputs (Ebrahimi et al., 2024), while studies on convincing evidence evaluation suggest that counter-stereotypical reasoning in prompts enhances fairness (Wan et al., 2024).
+Cognitive Mechanisms of Dual Process Theory.
 Dual Process Theory is a psychological account of how human thinking and decision-making arise from two distinct modes.
-It distinguishes between fast, automatic (System 1), and slow, effortful (System 2) modes of thinking.
 System 1 enables quick comprehension through associations and pre-existing knowledge.
 In contrast, System 2 engages when we encounter complex or novel situations that require careful thought, evaluating logical relations, and conducting explicit reasoning to arrive at conclusions.
-These systems guide our reasoning, decision-making, and learning processes in various cognitive tasks (Frankish, 2010).
-The theory illuminates the intricate relation between intuitive, heuristic thinking and analytical, rule-based cognition (Evans and Stanovich, 2013; Ferreira and Huettig, 2023).
-Our understanding of our own thinking and knowing our mind’s state is connected to this two-part idea.
-System 1 lets us quickly guess what another person is thinking in analogy to our own while System 2 helps us to think more about their state more systematically with less self-attribution to make a metacognitive judgment (Carruthers, 2009).
-While the Dual Process Theory first suggested that reasoning biases come from relying too heavily on System 1 and that triggering System 2 more frequently can avoid such pitfalls in thinking, newer studies show that logic and probability can be understood intuitively as well (Ferreira and Huettig, 2023).
+These systems guide our reasoning, decision-making, and learning processes in various cognitive tasks (Frankish, 2010; Evans and Stanovich, 2013; Ferreira and Huettig, 2023; Nighojkar et al., 2025).
+While the Dual Process Theory first suggested that reasoning biases come from relying too heavily on System 1 and that triggering System 2 more frequently can avoid such pitfalls in thinking, newer studies show that logic and probability can be understood intuitively as well (Ferreira and Huettig, 2023; Carruthers, 2009).
 Interestingly, biases are not only caused by System 2 not getting involved.
 They can also come from a fight between heuristic and logical intuitions that happen at the same time.
-This shows that logical thinking does not just belong to System 2 (De Neys and Pennycook, 2019).
-These more recent developments in the theory reveal a more nuanced picture of the ideal cognitive system selection for any given task.
-Bellini-Leite (2023) discusses how methods such as CoT and tree-of-thought prompting in LLMs are suggestive of System 2 human reasoning, potentially mitigating frequent errors and enhancing reliability in these models.
-Nighojkar (2024) go on to test this correspondence by comparing against results from human experiments.
-He finds that CoT prompting does not simply mimic System 2.
-Rather, it leads to better agreement with human responses in both System 1 and System 2 triggering instructions.
+Bellini-Leite (2023) highlights how CoT and tree-of-thought prompting align with System 2 reasoning, reducing errors and improving model reliability.
+Nighojkar (2024) tests this by comparing LLM outputs to human responses, finding that CoT prompting enhances agreement with both System 1 and System 2 reasoning rather than merely mimicking System 2.
+The Role of Personas in LLM.
 Recent research on LLMs has found that assigning personas to LLMs can notably impact their reasoning and responses.
 Beck et al.
-(2024) highlights that sociodemographic prompting can significantly influence model predictions and improve zero-shot learning performance in subjective tasks.
+(2024); de Araujo and Roth (2024); Kamruzzaman et al.
+(2024a) highlights that sociodemographic prompting can significantly influence model predictions and improve zero-shot learning performance in subjective tasks.
 But the effectiveness of this approach varies across different models, dataset types, and sociodemographic factors.
-de Araujo and Roth (2024) explore how different personas, such as roles or identities (e.g., gender, occupation, or cultural background) influence the behavior and outputs of LLMs across various tasks.
-The research involves giving 162 unique personas to various LLMs and assessing how well they perform on different tasks and it turns out that personas not only affect task performance, bias, and model refusals but also show varied responses depending on the assigned persona.
 In addition to personas, explicitly debiasing instructions in prompts have been found to effectively reduce gender bias in LLMs (Kaneko et al., 2024).
+While previous studies have explored dual process theory primarily in reasoning contexts (e.g., mathematical problem-solving), none have explicitly focused on reducing social bias.
+In this study, we investigate the intersection of dual process theory, persona, and social bias in LLMs.
+By integrating System 1 and System 2 with persona, we propose a distinct approach to reduce social biases in LLMs.
 3 Dataset
-We use two bias detection datasets to measure the behavior of the LLMs, StereoSet (Nadeem et al., 2020) and Kamruzzaman et al.’s (Kamruzzaman et al., 2024) bias detection dataset.
+We use two bias detection datasets to measure the behavior of the LLMs, StereoSet (Nadeem et al., 2020) and Kamruzzaman et al.’s (Kamruzzaman et al., 2024b) bias detection dataset.
 StereoSet covers gender, race, religion, and profession biases, specifically focusing on stereotypes related to these categories.
@@
 4 Experimental Setup
-We evaluate five major language models in this paper: 1) GPT-4 (Achiam et al., 2023), using the GPT-4 checkpoint on the OpenAI API; [^2] Due to the cost of running GPT-4, we use a sample of 2,100 entries from the GenAssocBias dataset to run GPT-4.
+We evaluate five major language models in this paper: (1) GPT-4222Due to the cost of running GPT-4, we use a sample of 2,100 entries from the GenAssocBias dataset to run GPT-4.
 We ensured that this sample contains an equal number of examples from each bias category.
-2) GPT-3.5 (Brown et al., 2020), using the GPT-3.5-TURBO-INSTRUCT checkpoint on the OpenAI API; 3) Llama-2-13B (Touvron et al., 2023), using the TheBloke/Llama-2-13B-chat-GGUF [^3] We use 4-bit quantized Llama 2 and Mistral models due to limited local computing resources.
-checkpoint on Huggingface; 4) Mistral-7B (Jiang et al., 2023), using the TheBloke/Mistral-7B-Instruct-v0.1-GGUF checkpoint on Huggingface; 5) Gemini (Team et al., 2023), using gemini-1.0-pro-latest on the Google Gemini API.
+(2) GPT-4o-mini (3) Llama3.3-70B (4) Mistral-7B (5) Gemma3-27B.
+Details about model selection see Appendix A.
 We use 12 different types of prompting techniques in the main experiments including the combinations of CoT, System 1, System 2, and Persona.
 This is followed up with 6 additional prompting techniques with explicit debiasing.
+We run each of the prompting techniques three times and report our results averaged across all 3 runs.
 Type Prompt
@@
 Given the context below, …
-HP + CoT Adopt the identity of a person who answers questions by thinking them “step-by-step".
+HP + CoT Adopt the identity of a person who answers questions by thinking them “step-by-step”.
 Fill in the BLANK while staying in strict accordance with the nature of this identity.
@@
 Their answers are effortful and reliable.
-They answer questions by thinking about them “step-by-step".
+They answer questions by thinking about them “step-by-step”.
 Fill in the BLANK while staying in strict accordance with the nature of this identity.
@@
 Given the context below, …
-MP + CoT Adopt the identity of a machine that answers questions by analyzing them “step-by-step".
+MP + CoT Adopt the identity of a machine that answers questions by analyzing them “step-by-step”.
 Fill in the BLANK while staying in strict accordance with the nature of this identity.
@@
 Its answers are effortful and reliable.
-It answers questions by analyzing them “step-by-step".
+It answers questions by analyzing them “step-by-step”.
 Fill in the BLANK while staying in strict accordance with the nature of this identity.
@@
 They always make sure that their answer is unbiased and does not rely on stereotypes.
-They answer questions by thinking about them “step-by-step".
+They answer questions by thinking about them “step-by-step”.
 Fill …
@@
 For our CoT prompting, we use the zero-shot variant which includes the phrase “Let’s think step-by-step” in the prompt (Kaneko et al., 2024).
-This method encourages the LLM to unfold its thinking process in a structured and detailed manner.
-This is the simplest variant of the prompting method often compared to System 2 reasoning.
+This method encourages the LLM to unfold its thinking process in a structured and detailed manner and has been shown to improve LLM capabilities in reasoning tasks.
 4.3 System 1 and System 2 Prompts
-System 1 prompting frames the context so that LLMs generate text reflecting the quick decision-making process in human cognition known as System 1.
+System 1 prompting frames the context so that LLMs generate text resembling [^3] We are deliberately not testing whether LLMs’ processing are subject to System 1 and System 2 processes themselves.
+Rather, the prompts place the textual context where the text generated by the LLMs will mimic text produced by people using System 1 or System 2 processes.
+the quick decision-making process in human cognition known as System 1.
 The System 1 prompt is “You answer questions quickly.
@@
 5 Results & Analysis
+We present our main results in terms of stereotypical engagement/response rates, indicating the percentage (%) of responses that aligned with stereotypical judgments.
+The ethical stance of this paper is that stereotypical judgments are a form of representational harm and our goal is to minimize such generations in LLMs.
+Stereotyping reduces individuals into beliefs about the groups that they are members of, which acts as a form of dehumanization, perpetuates existing inequalities, and marginalizes minorities.
+Overall Prompting Effects.
+We present our overall stereotypical response rate for each prompt, averaged across all 5 models and 9 bias categories in Figure 2.
+Figure 2 shows that on average a Human Persona with System 2 and CoT prompting best reduces social bias in LLMs.
+We also see that on average the standard prompting results is more stereotypical than other prompting techniques.
+We also observe that System 1 is more stereotypical than System 2.
 Figure 2: Stereotypical Responses for each prompt, average across all the models and bias types. Here, MP stands for Machine Persona, HP stands for Human Persona.
@@
 Here, MP stands for Machine Persona, HP stands for Human Persona.
-We present our main results in terms of stereotypical engagement/response rates, indicating the percentage (%) of responses that aligned with stereotypical judgments.
-Overall Prompting Effects.
-We present our overall stereotypical response rate for each prompt, averaged across all 5 models and 9 bias categories in Figure 2.
-Figure 2 shows that on average a Human Persona with System 2 prompting best reduces social bias in LLMs.
-We also see that on average the System 1 prompting is more stereotypical than other prompting techniques.
-This aligns with the behavior we would expect from dual process theory assuming that LLMs roughly model human cognition and that our System 1 and System 2 prompts appropriately trigger similar biases to the corresponding human cognitive systems.
-A surprising result is that CoT prompting does not show any reduction in bias.
-In fact, for every available minimal pair in prompts we consider (Standard vs.
-CoT, HP System 2 vs.
-HP System 2 + CoT, and MP System 2 vs.
-MP System 2 + CoT), CoT leads to an increase in stereotypical responses.
 Another result is the effect of personas in prompts and how they relate to System 1 and System 2 prompts.
 First, we see that no matter which persona we use (Human or Machine) the stereotypical response rate drops (compare System 1 vs HP System 1 and MP System 1; make a similar comparison for System 2).
-This suggests that having an LLM model as a separate entity (human or machine) leads to less socially biased outputs.
-When System 1 and System 2 prompts are combined with a human persona, their effects on social bias are amplified.
-The difference between the System 1 and System 2 responses is greater with the Human Persona + System 2 prompts having the least stereotypical responses overall.
-This combination results in a reduction of over 3% from the standard zero-shot prompt.
+This suggests that having an LLM model a separate entity (human or machine) leads to less socially biased outputs.
+When System 1 and System 2 prompts are combined with a human persona, the benefits of the prompts on social bias are amplified.
+The difference between the System 1 and System 2 responses is greater with the Human Persona.
+The Human Persona + System 2 + CoT prompts has the least stereotypical responses overall, with a reduction of around 5% from the standard zero-shot prompt.
 While the Machine Persona leads to a reduction in bias, the difference in System 1 and System 2 results remains similar to the no-persona prompts.
-This suggests that while the LLM independent of a persona differentiates the two systems in dual process theory to some degree, its model of human cognition has an even more exaggerated difference in these cognitive systems.
+This suggests that while the LLM’s generations differentiate the two systems in dual process theory to some degree independent of a persona, the LLM’s modeling of human-like generations has an even more exaggerated difference relative to these cognitive systems.
 5.1 Debiasing Prompt Follow-up
@@
 (2024).
-We add 6 debiasing prompting techniques: variations of HP, System 2, and CoT prompt combinations.
+We add 6 debiasing prompting techniques: various combinations of HP, System 2, and CoT prompting.
 The exact debasing prompts are shown in Table 2.
 Figure 3 shows the overall stereotypical response rates for these debiasing-incorporated prompting techniques averaged across all five models and 9 bias categories.
-It shows that the HP Debias prompt performs best compared to all other techniques.
+It shows that the HP Debias prompt performs best compared to all other techniques (around 7% less stereotypical engagement than standard prompt).
 Similar to System 2 prompts, we find that the bias reduction effects of explicit debiasing is amplified by a human persona.
-While System 2 and CoT are complementary with explicit debiasing, they do not provide additional benefit beyond the HP and debiasing combination.
-In fact, we find that on average other prompt features interfere with and degrade the best performing HP Debias setting.
-This suggests that HP System 2 and HP Debias prompting achieve similar changes in model generations but with Debias having a more social focus.
-This is not surprising when we look at the prompts.
-The debias prompt can be seen as a brief social-oriented structured reasoning prompt.
+Although the HP Debias is the best performing techniques on average, the HP + System 2 + CoT + Debias technique also reduce social biases and difference between HP + Debias and HP + System 2 + CoT + Debias is very small (0.20%).
+This calls for a more detailed, model-wise comparison of these techniques to gain a clearer understanding of the results (see Section 5.3).
+Type Age Beauty Beauty-P. Insti. Nation. Gender Prof. Race Religion Avg.
+GPT-4o-mini
+Standard Prompt 51.84 73.91 54.89 63.77 50.42 77.60 77.62 59.53 59.74 63.26
+HP + Debias -8.04* $\downarrow$ -11.31* $\downarrow$ -8.85* $\downarrow$ -6.93* $\downarrow$ -4.65* $\downarrow$ -10.66* $\downarrow$ -13.93* $\downarrow$ -16.47* $\downarrow$ -10.39* $\downarrow$ -10.14 $\downarrow$
+HP+System 2+CoT+Debias -5.53 $\downarrow$ -4.89* $\downarrow$ -4.26* $\downarrow$ -5.04 $\downarrow$ -1.71 $\downarrow$ -8.02* $\downarrow$ -16.40* $\downarrow$ -13.54* $\downarrow$ -5.69* $\downarrow$ -7.23 $\downarrow$
+HP + System 2 -4.12 $\downarrow$ -1.94 $\downarrow$ -0.42 $\downarrow$ -0.23 $\downarrow$ +0.54 $\uparrow$ -3.87* $\downarrow$ -6.65* $\downarrow$ -3.71* $\downarrow$ -0.84 $\downarrow$ -2.36 $\downarrow$
+HP + System 2 + Debias -5.92* $\downarrow$ -7.53* $\downarrow$ -5.17* $\downarrow$ -4.53* $\downarrow$ -2.45 $\downarrow$ -4.12* $\downarrow$ -14.84* $\downarrow$ -12.53* $\downarrow$ -9.74* $\downarrow$ -7.43 $\downarrow$
+Gemma3-27B
+Standard Prompt 42.06 72.45 48.08 27.59 40.50 69.46 63.72 60.73 57.14 53.53
+HP + Debias -1.10 $\downarrow$ -6.37* $\downarrow$ +1.15 $\uparrow$ -1.06 $\downarrow$ -1.03 $\downarrow$ -0.86 $\downarrow$ -2.10 $\downarrow$ -12.59* $\downarrow$ -2.97 $\downarrow$ -3.00 $\downarrow$
+HP+System 2+CoT+Debias -2.26 $\downarrow$ -6.38* $\downarrow$ -1.00 $\downarrow$ -1.74 $\downarrow$ -2.66 $\downarrow$ -2.66 $\downarrow$ -0.44 $\downarrow$ -11.68* $\downarrow$ -3.72 $\downarrow$ -3.62 $\downarrow$
+HP + System 2 -7.12* $\downarrow$ -3.82* $\downarrow$ -1.07 $\downarrow$ -2.22 $\downarrow$ -6.67* $\downarrow$ -6.65 $\downarrow$ -4.88* $\downarrow$ -6.64* $\downarrow$ -0.98 $\downarrow$ -3.62 $\downarrow$
+HP + System 2 + Debias -4.33* $\downarrow$ -6.77* $\downarrow$ -2.43 $\downarrow$ -2.04 $\downarrow$ -3.78* $\downarrow$ -3.07* $\downarrow$ -1.88 $\downarrow$ -11.06* $\downarrow$ -0.20 $\downarrow$ -3.96 $\downarrow$
+Llama3.3-70B
+Standard Prompt 40.85 75.15 45.21 36.57 47.37 74.68 67.89 58.69 50.79 55.24
+HP + Debias -11.10* $\downarrow$ -28.40* $\downarrow$ -15.53* $\downarrow$ -5.14 $\downarrow$ -8.55* $\downarrow$ -12.65* $\downarrow$ -16.78* $\downarrow$ -19.73* $\downarrow$ -20.36* $\downarrow$ -15.24 $\downarrow$
+HP+System 2+CoT+Debias -12.27* $\downarrow$ -32.81* $\downarrow$ -24.43* $\downarrow$ -5.57* $\downarrow$ -12.40* $\downarrow$ -19.26* $\downarrow$ -17.82* $\downarrow$ -18.71* $\downarrow$ -20.94* $\downarrow$ -18.24 $\downarrow$
+HP + System 2 -8.40* $\downarrow$ -22.49* $\downarrow$ -20.58* $\downarrow$ -2.46 $\downarrow$ -5.33* $\downarrow$ -10.10* $\downarrow$ -12.29* $\downarrow$ -10.32* $\downarrow$ -11.40* $\downarrow$ -11.48 $\downarrow$
+HP + System 2 + Debias -10.12* $\downarrow$ -32.06* $\downarrow$ -22.62* $\downarrow$ -4.85 $\downarrow$ -11.78* $\downarrow$ -19.64* $\downarrow$ -19.04* $\downarrow$ -18.88* $\downarrow$ -15.50* $\downarrow$ -17.16 $\downarrow$
+Table 3: Comparison of changes in stereotypical response rates for selected prompting techniques across GPT-4o-mini, Gemma3, and Llama3.3.
+Results reported are compared to the Standard Prompt (increased from the Standard prompt: $\uparrow$ in red, decrease: $\downarrow$ in green).
+Least stereotypical responses of each bias category-model are bolded.
+Avg.
+is the macro average of each prompting technique.
+* denotes statistically significant results (except Avg.
+column) compared to the standard prompt using Kendall’s $\tau$ test Kendall (1938).
 5.2 Model- and Bias-specific Prompting Effects
@@
 All of the standard prompting results alongside the best performing prompting technique results for each bias category and model combination are presented in Figure 4.
-Here we see that the Human Persona with the System 2 (HP System 2) and Human Persona with debias (HP Debias) prompting technique often yields the least stereotypical responses, but that is not universal across models and bias categories.
-HP Debias outperforms all other prompting techniques in 14 cases.
-Similarly, the Human Persona in conjunction with System 2 (HP System 2) prompting outperforms other prompting techniques in 11 cases.
-Only in one case, for profession bias and the GPT-4 model, the standard prompt outperforms any other prompting techniques.
-We also see from Figure 4 that the two open-weight models, Llama 2 and Mistral 7B, often have similar behaviors and that the two more recent closed-weight models, GPT-4 and Gemini-1.0 have similar behaviors.
-GPT-4 stands out in having the most cases where prompting variants make minimal improvement from the standard zero-shot prompt.
-This suggests that OpenAI has done some engineering on this front.
-Though it is unclear whether this is behind-the-scenes prompt modifications, analogous instruction finetuning of the model, or some other method to make the model robust to prompt variants.
-Next, we focus on each bias category.
-Ageism.
-We see no consistent prompt setting that performs best on ageism. Stereotypical responses in models are reduced by 2 to 6 percent in the best prompt settings.
+Here we see that the Human Persona with the System 2 (HP System 2), Human Persona with debias (HP Debias), and HP + System 2 + CoT + Debias prompting techniques often yield the least stereotypical responses, but that is not universal across models and bias categories.
+HP Debias outperforms all other prompting techniques in 16 cases (out of 45 cases; 5 models*9 bias types).
+Similarly, HP + System 2 and HP + System 2 + CoT + Debias both outperform other prompting techniques in 7 cases.
+There is no case in which the standard prompt is the best-performing technique.
+Ageism. We see no consistent prompt setting that performs best on ageism. Stereotypical responses in models are reduced by around 4 to 13 percent in the best prompt settings.
 Beauty.
-Prompt variants in our experiments show substantial improvements on beauty bias across all considered models—up to 19 percent reduction in stereotypical responses in Mistral-7B using the HP System 2 prompt.
-The remaining 4 models also show major improvements in beauty bias, all using the HP Debias prompt.
+Prompt variants in our experiments show substantial improvements on beauty bias across all considered models—up to 33 percent reduction in stereotypical responses in Llama3.3 using the HP + System 2 + CoT + Debias prompt.
+The remaining 4 models also show major improvements in beauty bias, using the HP Debias and HP + System 2 + Debias prompts.
 Beauty in Profession.
-Gemini-1.0 shows a 13 percent reduction in stereotypical responses for beauty in profession bias using System 2 Debias prompting.
-The best prompt setting is inconsistent for this bias category, with HP System 2 showing the largest bias reduction for GPT-4 and Llama 2 while HP Debias and HP System 2 Debias results in the largest bias reduction in GPT-3.5 and Mistral-7B, respectively.
+Llama3.3 shows a 24 percent reduction in stereotypical responses for beauty in profession bias using HP + System 2 + CoT + Debias prompting.
+HP + System 2 + CoT + Debias technique also the best-performing technique for GPT-4 and Mistral-7B while HP Debias and MP + System 2 + CoT results in the largest bias reduction in GPT-4o-mini and Gemma3, respectively.
 Gender.
 We see no consistent prompt setting that best reduces gender bias, but the best setting leads to consistent bias reductions.
-Interestingly, the open-weight models, Llama-2 and Mistral-7B show less stereotypical gender responses before explored prompt-based improvements than the other models after.
+Interestingly, among the open-weight models, Llama3.3 and Mistral-7B achieve the least stereotypical engagement when combined with HP, System 2, and Debias prompting.
+For the closed-source models, GPT-4 and GPT-4o-min produce the least stereotypical responses with HP Debias prompting.
 Institutional.
@@
 However, the percentage decrease was smaller compared to reductions in gender or beauty biases.
-With Llama 2, we achieved about a 5 percent improvement when using HP System 1 for prompting.
+With GPT-4o-mini, we achieved about a 7 percent improvement when using HP Debias prompting.
 Nationality.
 Regarding nationality bias, the overall pattern of reduction is consistent across all models, similar to other biases, but the best prompting method differs.
-The combination of HP System 2 with CoT delivers the best performance for GPT-3.5 and Mistral 7B.
 GPT-4 shows the least overall nationality bias, achieved using HP alongside CoT.
 Profession.
-We achieved up to an 11 percent reduction in bias for the profession category.
-Here we see the single bias-model combination where standard prompting performed best in GPT-4.
+We achieved up to a 19 percent reduction in bias for the profession.
 The other models all show substantial improvements.
-GPT-3.5 and Gemini-1.0, System 2 prompts yielded the best results.
-For Mistral-7B, Hp System 2 and for Llama 2, HP Debias were the best.
+Mistral-7B and Gemma3, HP + System 2 prompts yielded the best results.
+For GPT-4o-mini, System 2 + Debias and for GPT-4, HP Debias were the best.
 Race.
-We observe a reduction in racial biases across all models, although the decrease is relatively small for GPT-4, akin to that observed with standard prompting techniques.
-In contrast, the Mistral 7B model using HP System 2 with CoT prompting shows a bias reduction of approximately 9 percent.
-The best-performing prompting technique for race is HP Debias for GPT-3.5 which reduces around 17 percent of stereotypical engagement.
-Here HP Debias was the best prompting technique in three of the five models.
+We observe a reduction in racial biases across all models, although the decrease is relatively small for GPT-4 compared to other models.
+HP Debias is the best-performing technique for four models with Llama3.3 shows a bias reduction of approximately 20 percent.
+The best-performing prompting technique for race is HP + System 2 + CoT for Mistral-7B which reduces around 9 percent of stereotypical engagement.
 Religion.
-We achieved a reduction in religious bias by up to 13 percent.
-Additionally, we observed reductions across all models, although the decrease in the GPT-4 model was relatively minor.
+We achieved a reduction in religious bias by up to 21 percent.
+Additionally, we observed reductions across all models, although the decrease in the GPT-4 model was relatively small.
 Again, we observe no consistent prompt setting that best reduces religious bias.
-6 Does CoT Prompting Best Model System 2?
+5.3 Model-wise Discussion and Suggestions
+From the previous Section 5.2, we can see that in most cases no single prompting technique is consistently reduce biases across bias-model category combinations, that require more fine grained analysis of results in individual model-wise.
+Table 3 presents a few selected prompting techniques’ results for GPT-4o-mini, Gemma3, and Llama3.3.
+Statistical testing results for a few selected cases are listed in Table 9 in Appendix C.
+For the complete results of each model and a more detailed discussion, see Appendix B.
+Most of the selected prompting techniques reduce biases in GPT-4o-mini, with HP Debias being the best-performing technique across 8 bias categories.
+From Table 3, we observe that the HP Debias technique performs best for GPT-4o-mini compared to other methods.
+On average, models tend to perform better when explicit debiasing instructions are included than when they are not.
+Therefore, users who lack deep knowledge of bias reduction techniques and wish to use closed-source GPT models can consider using the HP Debias technique as an effective approach.
+For Gemma3, most of the selected prompting techniques reduce biases, although the reduction is smaller compared to other models.
+However, each technique reduces biases to a similar extent, as reflected in the similar average scores.
+We observe a smaller reduction in bias with the Gemma3 model.
+For Gemma3, the HP Debias technique generally does not perform well, except in the race bias category.
+However, the HP + System 2 and HP + System 2 + CoT + Debias techniques perform better in most cases.
+Therefore, those considering the use of Gemma3 may find these two methods more effective.
+The bias reduction rate for Llama3.3 is higher than any other models and HP + System 2 + CoT + Debias is the best-performing technique.
+From Table 3, we observe that most techniques substantially reduce biases (larger reduction), with this reduction being consistent across the selected methods.
+Therefore, users looking to mitigate biases in open-source models without fine-tuning can apply these debiasing techniques effectively with Llama3.3.
+We also observe that the GPT models exhibit similar behavior regarding the best-performing techniques, with HP Debias emerging as the most effective across most bias categories for both models in 15 cases (see Table 4 and Table 5).
+So, out of 16 best performing HP Debias cases, 15 are from GPT models.
+We also observe that GPT-4 demonstrates a smaller overall reduction in bias compared to GPT-4o-mini.
+6 Conclusion
+Our study has contributed to the understanding and reduction of social biases in LLMs through prompting techniques inspired by dual process theory.
+By testing the effects of System 1 and System 2 textual contexts, as well as the incorporation of human-like personas and debiasing prompts, our research not only clarifies the relationship between dual process theory and the generative patterns of LLMs but also demonstrates practical methods for reducing biases in LLM generations.
+Our findings reveal that System 2 prompts, particularly when combined with a Human Persona, consistently reduce stereotypical judgments across various social bias categories.
+Biases were further reduced when using a debiasing prompt, which can be seen as a social bias-focused System 2 prompt, along with a human persona.
+These prompt variations are simple prefixes to the main instructions, making this technique straightforward to incorporate in most LLM use cases.
+This indicates a profound potential for combining analysis-leaning instructions and personalized prompting to enhance the ethical performance of LLMs.
+Furthermore, our use of different models and bias datasets ensures our results are robust and applicable across different contexts.
+7 Limitations
+Limitations of the System 1 and System 2 analogies as they extend to LLMs.
+The analogy of dual process theory over LLMs is far from straightforward.
+LLMs process information in a fundamentally different manner than humans, so any analogy must grapple with this.
+This could be accomplished by forcing the LLM to have similar (or analogous) processing limitations that humans undergo in certain scenarios (e.g., time sensitivity $\leftrightarrow$ number of processing cycles).
+Alternatively, as we do here, we only discuss dual process theory’s effects as it is approximated by LLMs via textual training.
+Since people generate the text that LLMs are trained on, the effects of dual process theory will be approximated by LLMs in similar textual contexts regardless of the underlying processing method of the LLM.
+We do not investigate the relationship between dual process theory and the internal process of LLMs, only the generative patterns relative to textual context.
+Our study is designed to investigate the behavior of System 1 and System 2-based prompting.
+The conditions under which people use System 1 and System 2 reasoning have been well-studied and are reflected in our prompts.
+We do not aim to measure if the LLM actually processes information in a way that reflects System 1 or System 2 reasoning as these systems are inherent to human cognition.
+We use the prompts merely to access the LLM’s approximation of these cognitive systems under related textual contexts.
+Desired generative distributions.
+In certain scenarios, balancing stereotypical and anti-stereotypical statements may appear justifiable; however, the conditions that warrant such balancing are not explicitly delineated in the dataset we utilized.
+For example, while some stereotypes may be considered harmless on their own (only a problem when always assumed) or subject to cultural context, others, such as ‘Every single Muslim I ever met was clearly a terrorist’, are fundamentally harmful and must be avoided altogether.
+The inclusion of such extreme and offensive stereotypes underscores the need for a clear framework to distinguish between stereotypes that can be balanced and those that should be excluded entirely.
+We encourage future research to investigate methodologies for systematically categorizing and separating these stereotypes.
+This includes devising guidelines to identify harmful stereotypes, assessing their ethical implications, and determining appropriate responses that ensure both fairness and inclusivity without inadvertently reinforcing biased narratives.
+Scope of dataset evaluation.
+Our study evaluates LLM biases using two datasets, both in English and structured in a specific question-answering format.
+While this design ensures consistency in evaluation, it limits the generalizability of our findings to other data formats, such as open-ended responses, dialogue-based interactions, or real-world text corpora.
+Additionally, the exclusive use of English datasets prevents us from assessing whether the observed bias reduction effects extend to multilingual or culturally diverse contexts.
+Future research should explore a broader range of datasets, languages, and task types to determine the robustness of these findings.
+Trade-offs in model performance.
+While our study focuses on the effectiveness of bias reduction strategies, we do not assess whether the introduced prompts impact the overall performance of the model on the task itself.
+Factors such as fluency, coherence, and potential trade-offs in response quality remain unexamined.
+It is possible that certain debiasing techniques could inadvertently reduce model accuracy or alter response style in unintended ways.
+Future research should investigate these trade-offs to ensure that bias reduction does not come at the cost of overall task performance.
+Number of LLMs tested.
+In our paper, we tested five LLMs.
+However, due to resource constraints, we were unable to test other available models.
+Therefore, the results presented in this paper may not fully generalize to all LLMs.
+Acknowledgments
+This project was fully supported by the University of South Florida.
+Appendix A LLMs Selection
+We use five language models in this paper: 1) GPT-4 (Achiam et al., 2023), using the GPT-4 checkpoint on the OpenAI API; 2) GPT-4o-mini, using the GPT-4o-mini checkpoint on the OpenAI API; 3) Llama3.3-70B (Touvron et al., 2023), using Ollama [^4] https://ollama.com/ 4) Mistral-7B (Jiang et al., 2023), using Ollama; 5) Gemma3-27B (Team et al., 2024), using ollama.
+We selected LLMs to reflect common usage while balancing our research budget.
+We use a mix of commercial and open-weight systems.
+GPT remains the most common commercial LLM, and Llama, Gemma, and Mistral are popular open-weight LLMs that we could fit into our computing resources.
+All models are used with their default hyperparameter settings.
+Appendix B Model-wise results for all prompting techniques
+We presented our model wise results in Tables 4, 5, 6, 7 and 8
+B.1 Discussion of GPT-4o-mini Results
+Table 4 provides the results of GPT-4o-mini’s stereotypical biases across various prompting strategies.
+The results are presented in comparison to the Standard Prompt, highlighting reductions or increases in bias using different techniques such as CoT, System 1 and System 2 reasoning, and debiasing combinations.
+Observations on Standard Prompt and Bias Levels.
+The Standard Prompt demonstrates a baseline level of stereotypical responses across the categories of Age, Beauty, Gender, Institutional, Nationality, Profession, Race, and Religion.
+High values in Gender (77.60) and Profession (77.62) indicate areas where GPT-4o-mini is particularly prone to stereotyping.
+Other notable categories, such as Beauty (73.91) and Institution (63.77), also show substantial stereotypical bias levels.
+Effectiveness of Various Prompting Techniques.
+  • CoT: CoT slightly reduces biases across most categories, with the largest improvements in Religion (-3.16↓) and Institution (-2.62↓).
+However, the technique underperforms in reducing bias in Gender (+2.24↑), suggesting that while CoT aids reasoning, it does not universally reduce all bias types.
+  • System 1 and System 2:
+    System 1: While Beauty bias increases (+9.87↑), biases in categories like Institutional (-5.13↓) and Ageism (-5.95↓) decrease. However, the overall Avg. bias reduction is marginal (-0.26↓).
+    System 2: System 2 shows bias reduction across categories such as Profession (-7.48↓) and Religion (-6.49↓), with an average improvement of -2.68↓, making it more effective compared to System 1.
+  • HP and Combined Techniques:
+    HP + System 1/System 2: Adding human persona with System 1 or System 2 doesn’t really help GPT-4o-mini to reduce biases with HP + System 1 even increase the overall average bias (+0.46↑).
+  • HP with debiasing techniques yields best reductions across most categories, particularly Beauty (-11.31↓), Gender (-10.66↓), and Race (-16.47↓), achieving the best overall reduction (-10.14↓).
+  • System 2 with Debiasing: Combining System 2 with debiasing techniques yields substantial reductions across all categories, particularly Profession (-15.70↓), and Race (-11.09↓), achieving the overall reduction (-7.35↓).
+These results highlight System 2’s compatibility with debiasing methods for tackling deep-rooted biases.
+  • HP + System 2 + CoT + Debias: HP + System 2 + CoT + Debias achieves a substantial average improvement (-7.23↓) with notable reductions in Age (-5.53↓), Gender (-8.02↓), and Race (-13.54↓), confirming the effectiveness of combining approaches with explicit debiasing strategies.
+The results indicate that the human persona with debiasing techniques achieves the greatest reduction in stereotypical responses across most bias categories.
+Practitioners should favor this method for reducing biases in models like GPT-4o-mini.
+Failure cases.
+A closer inspection of Table 4 reveals that certain techniques can inadvertently increase bias in specific categories.
+For instance, System 2 fails for Beauty and Nationality, where biases increase by 5.20 and 2.92 respectively.
+HP + System 1 amplifies biases in Race and Religion by 2.05 and 3.27, while HP + CoT exhibits higher biases in Beauty-P, Gender, and Nationality.
+MP + CoT shows a considerable jump in Religion bias of 6.93.
+These cases underscore the complexity of mitigating biases across multiple dimensions, where reducing bias in one category can sometimes lead to regressions in others.
+Type Age Beauty Beauty-P. Gender Insti. Nation. Prof. Race Religion Avg.
+Standard Prompt 51.84 73.91 54.89 77.60 63.77 50.42 77.62 59.53 59.74 63.26
+CoT -0.42 $\downarrow$ -1.27 $\downarrow$ +0.56 $\uparrow$ +2.24 $\uparrow$ -2.62 $\downarrow$ +0.12 $\uparrow$ -0.82 $\downarrow$ -0.45 $\downarrow$ -3.16 $\downarrow$ -0.56 $\downarrow$
+System 1 -5.95 $\downarrow$ +9.87 $\uparrow$ +1.91 $\uparrow$ -0.89 $\downarrow$ -5.13 $\downarrow$ +2.32 $\uparrow$ -2.78 $\downarrow$ -0.38 $\downarrow$ -1.30 $\downarrow$ -0.26 $\downarrow$
+System 2 -6.05 $\downarrow$ +5.20 $\uparrow$ -1.24 $\downarrow$ -4.21 $\downarrow$ -4.35 $\downarrow$ +2.92 $\uparrow$ -7.48 $\downarrow$ -2.42 $\downarrow$ -6.49 $\downarrow$ -2.68 $\downarrow$
+HP + System 1 +1.02 $\uparrow$ +0.19 $\uparrow$ -0.40 $\downarrow$ +0.59 $\uparrow$ -1.54 $\downarrow$ -0.80 $\downarrow$ -0.19 $\downarrow$ +2.05 $\uparrow$ +3.27 $\uparrow$ +0.46 $\uparrow$
+HP + System 2 -4.12 $\downarrow$ -1.94 $\downarrow$ -0.42 $\downarrow$ -3.87 $\downarrow$ -0.23 $\downarrow$ +0.54 $\uparrow$ -6.65 $\downarrow$ -3.71 $\downarrow$ -0.84 $\downarrow$ -2.36 $\downarrow$
+HP + CoT -0.21 $\downarrow$ -0.65 $\downarrow$ +0.52 $\uparrow$ +0.63 $\uparrow$ -2.62 $\downarrow$ +0.04 $\uparrow$ -1.25 $\downarrow$ +0.88 $\uparrow$ -6.41 $\downarrow$ -0.90 $\downarrow$
+HP + System 2 + CoT -3.17 $\downarrow$ -2.35 $\downarrow$ -1.15 $\downarrow$ -6.96 $\downarrow$ -1.82 $\downarrow$ -0.22 $\downarrow$ -7.44 $\downarrow$ -3.82 $\downarrow$ -4.95 $\downarrow$ -3.54 $\downarrow$
+MP + System 1 -0.51 $\downarrow$ +0.16 $\uparrow$ -1.77 $\downarrow$ -0.31 $\downarrow$ -1.56 $\downarrow$ +0.40 $\uparrow$ +0.44 $\uparrow$ +3.16 $\uparrow$ -0.53 $\downarrow$ -0.06 $\downarrow$
+MP + System 2 -2.72 $\downarrow$ -0.20 $\downarrow$ -2.83 $\downarrow$ -4.16 $\downarrow$ -0.93 $\downarrow$ -0.28 $\downarrow$ -3.83 $\downarrow$ -1.73 $\downarrow$ -7.69 $\downarrow$ -2.71 $\downarrow$
+MP + CoT -1.44 $\downarrow$ +0.61 $\uparrow$ -1.83 $\downarrow$ -0.46 $\downarrow$ -2.15 $\downarrow$ +0.14 $\uparrow$ -0.29 $\downarrow$ +2.26 $\uparrow$ +6.93 $\uparrow$ +0.42 $\uparrow$
+MP + System 2 + CoT -4.31 $\downarrow$ -0.66 $\downarrow$ -1.39 $\downarrow$ -4.99 $\downarrow$ -1.41 $\downarrow$ +0.76 $\uparrow$ -4.73 $\downarrow$ -3.15 $\downarrow$ -2.59 $\downarrow$ -2.43 $\downarrow$
+Standard Prompt + Debias -4.05 $\downarrow$ -2.15 $\downarrow$ -3.28 $\downarrow$ -0.27 $\downarrow$ -2.69 $\downarrow$ -3.06 $\downarrow$ -4.90 $\downarrow$ -4.89 $\downarrow$ -6.49 $\downarrow$ -3.53 $\downarrow$
+HP + Debias -8.04 $\downarrow$ -11.31 $\downarrow$ -8.85 $\downarrow$ -10.66 $\downarrow$ -6.93 $\downarrow$ -4.65 $\downarrow$ -13.93 $\downarrow$ -16.47 $\downarrow$ -10.39 $\downarrow$ -10.14 $\downarrow$
+System 2 + Debias -4.73 $\downarrow$ -6.89 $\downarrow$ -5.42 $\downarrow$ -9.29 $\downarrow$ -2.93 $\downarrow$ -1.67 $\downarrow$ -15.70 $\downarrow$ -11.09 $\downarrow$ -8.39 $\downarrow$ -7.35 $\downarrow$
+HP + System 2 + Debias -5.92 $\downarrow$ -7.53 $\downarrow$ -5.17 $\downarrow$ -4.12 $\downarrow$ -4.53 $\downarrow$ -2.45 $\downarrow$ -14.84 $\downarrow$ -12.53 $\downarrow$ -9.74 $\downarrow$ -7.43 $\downarrow$
+CoT + Debias -6.14 $\downarrow$ -4.92 $\downarrow$ -4.39 $\downarrow$ -3.43 $\downarrow$ -5.43 $\downarrow$ -2.95 $\downarrow$ -10.54 $\downarrow$ -11.96 $\downarrow$ -9.74 $\downarrow$ -6.61 $\downarrow$
+HP + System 2 + CoT + Debias -5.53 $\downarrow$ -4.89 $\downarrow$ -4.26 $\downarrow$ -8.02 $\downarrow$ -5.04 $\downarrow$ -1.71 $\downarrow$ -16.40 $\downarrow$ -13.54 $\downarrow$ -5.69 $\downarrow$ -7.23 $\downarrow$
+Table 4: Results for GPT-4o-mini.
+Results reported compared to Standard Prompt (increased from the Standard prompt: $\uparrow$ in red, decrease: $\downarrow$ in green).
+Least stereotypical responses of each bias category are bolded.
+Avg.
+is the macro average of each prompting technique.
+B.2 Discussion of GPT-4 Results
+Table 5 reports the observed stereotypical biases of GPT-4 under a variety of prompting techniques.
+Each approach is compared with the Standard Prompt baseline, enabling a clear comparison of how Syetme 2, human persona (HP) integration, debiasing, and their combinations influence the model’s responses across multiple bias categories.
+Observations on Standard Prompt and Baseline Biases.
+Under the Standard Prompt, GPT-4 exhibits notable stereotypical engagement in certain domains.
+High baseline values in Gender (77.65) and Profession (68.90) suggest that these domains remain particularly challenging, while Beauty (69.25) also shows elevated bias levels.
+In contrast, categories like Institutional (29.36) and Nationality (27.94) start from relatively lower bias scores.
+Impact of Human Persona.
+CoT prompting technique produces mixed results, slightly decreasing biases in some categories but increasing them in others.
+System 1 vs.
+System 2 reveals nuanced patterns.
+System 1 strategies produce incremental changes, sometimes increasing bias in domains like Beauty, whereas System 2 offering slight reductions in Age and other categories but occasionally increases bias elsewhere.
+Introducing a Human Persona (HP) into the prompt also yields complex effects.
+While HP combined with System 2 reduces biases in categories such as Gender (-4.53↓) and Beauty (-4.46↓), integrating HP with System 1 results in less consistent improvements, sometimes increasing the overall average bias.
+Efficacy of Debiasing Techniques and Their Combinations.
+Explicit debiasing methods, whether employed alone or coupled with System 2 or CoT, often produce substantial reductions.
+For instance, HP + Debias stands out for achieving broad improvements across several challenging categories, including significant reductions in Age (-11.77↓), Beauty (-10.40↓), and Gender (-11.44↓).
+Similarly, CoT + Debias, System 2 + Debias, and especially HP + System 2 + CoT + Debias consistently lower biases in traditionally difficult domains such as Race and Profession.
+However, these gains sometimes accompany trade-offs, as certain categories (like Institution or Nationality) may see increased bias levels even within broadly improved configurations.
+Failure cases.
+A closer examination of Table 5 reveals multiple instances where various methods inadvertently produce greater bias than the standard prompt.
+For example, CoT shows increases in Beauty-P, Institution, Nationality, Profession, Race, and Religion, while System 1 leads to higher biases for Beauty, Institution, Profession, and Race.
+HP + System 1 amplifies bias in Beauty, Gender, Nationality, Profession, and Race, whereas System 2 also raises bias for Beauty, Gender, Institution, Nationality, Profession, and Religion.
+Different prompt combinations can exhibit similar failures, such as MP + CoT or MP + System 2 inflating biases in Beauty, Beauty-P, Profession, Race, or Religion.
+Even when explicit debias prompts are used, several methods still substantially increase bias in Institution or Nationality; for instance, System 2 + Debias, HP + Debias, and CoT + Debias all show large jumps in those categories.
+Type Age Beauty Beauty-P. Gender Insti. Nation. Prof. Race Religion Avg.
+Standard Prompt 50.74 69.25 46.79 77.65 29.36 27.94 68.90 54.16 53.85 53.18
+CoT -1.97 $\downarrow$ -1.77 $\downarrow$ +1.62 $\uparrow$ -1.37 $\downarrow$ +0.86 $\uparrow$ +0.84 $\uparrow$ +2.94 $\uparrow$ +0.12 $\uparrow$ +1.28 $\uparrow$ +0.29 $\uparrow$
+System 1 -0.24 $\downarrow$ +0.95 $\uparrow$ -2.87 $\downarrow$ -0.57 $\downarrow$ +1.80 $\uparrow$ -0.38 $\downarrow$ +4.24 $\uparrow$ +1.11 $\uparrow$ -0.60 $\downarrow$ +0.38 $\uparrow$
+System 2 -3.09 $\downarrow$ +1.53 $\uparrow$ -0.14 $\downarrow$ +0.39 $\uparrow$ +1.19 $\uparrow$ +0.46 $\uparrow$ +2.53 $\uparrow$ -0.20 $\downarrow$ +1.99 $\uparrow$ +0.52 $\uparrow$
+HP + System 1 -4.55 $\downarrow$ +2.95 $\uparrow$ -0.81 $\downarrow$ +0.22 $\uparrow$ -0.34 $\downarrow$ +0.01 $\uparrow$ +3.84 $\uparrow$ +0.36 $\uparrow$ -1.29 $\downarrow$ +0.04 $\uparrow$
+HP + System 2 -3.16 $\downarrow$ -4.46 $\downarrow$ +2.20 $\uparrow$ -4.53 $\downarrow$ +0.62 $\uparrow$ +0.87 $\uparrow$ +1.73 $\uparrow$ +0.30 $\uparrow$ -1.22 $\downarrow$ -0.85 $\downarrow$
+HP + CoT -2.88 $\downarrow$ +1.02 $\uparrow$ +1.19 $\uparrow$ -2.36 $\downarrow$ 0.00 -0.95 $\downarrow$ +4.05 $\uparrow$ +0.74 $\uparrow$ -1.90 $\downarrow$ -0.12 $\downarrow$
+HP + System 2 + CoT -4.77 $\downarrow$ -0.13 $\downarrow$ +0.30 $\uparrow$ -2.85 $\downarrow$ +0.95 $\uparrow$ +0.25 $\uparrow$ +1.70 $\uparrow$ +1.26 $\uparrow$ 0.00 -0.36 $\downarrow$
+MP + System 1 -2.44 $\downarrow$ +3.68 $\uparrow$ +0.98 $\uparrow$ -3.34 $\downarrow$ +0.07 $\uparrow$ +0.18 $\uparrow$ +2.10 $\uparrow$ +0.88 $\uparrow$ -0.60 $\downarrow$ +0.17 $\uparrow$
+MP + System 2 -5.97 $\downarrow$ +1.31 $\uparrow$ +2.08 $\uparrow$ -1.18 $\downarrow$ +0.45 $\uparrow$ -0.47 $\downarrow$ +2.81 $\uparrow$ -0.30 $\downarrow$ +1.99 $\uparrow$ +0.08 $\uparrow$
+MP + CoT -3.68 $\downarrow$ +1.93 $\uparrow$ +0.16 $\uparrow$ -2.75 $\downarrow$ -0.86 $\downarrow$ -0.47 $\downarrow$ +4.73 $\uparrow$ +0.76 $\uparrow$ -0.52 $\downarrow$ -0.08 $\downarrow$
+MP + System 2 + CoT -5.02 $\downarrow$ -0.15 $\downarrow$ +2.17 $\uparrow$ -0.18 $\downarrow$ -0.20 $\downarrow$ +0.25 $\uparrow$ +3.56 $\uparrow$ +0.67 $\uparrow$ -1.90 $\downarrow$ -0.09 $\downarrow$
+Standard Prompt + Debias -2.99 $\downarrow$ -1.82 $\downarrow$ -3.19 $\downarrow$ -2.42 $\downarrow$ +23.61 $\uparrow$ +18.14 $\uparrow$ -3.13 $\downarrow$ +0.39 $\uparrow$ 0.00 +3.18 $\uparrow$
+HP + Debias -11.77 $\downarrow$ -10.40 $\downarrow$ -4.14 $\downarrow$ -11.44 $\downarrow$ +19.49 $\uparrow$ +16.85 $\uparrow$ -8.81 $\downarrow$ -6.69 $\downarrow$ -2.57 $\downarrow$ -2.16 $\downarrow$
+System 2 + Debias -9.62 $\downarrow$ -5.17 $\downarrow$ -4.74 $\downarrow$ -7.06 $\downarrow$ +24.56 $\uparrow$ +12.17 $\uparrow$ -6.80 $\downarrow$ -0.21 $\downarrow$ -1.29 $\downarrow$ +0.21 $\uparrow$
+HP + System 2 + Debias -10.36 $\downarrow$ -3.99 $\downarrow$ -3.57 $\downarrow$ -6.29 $\downarrow$ +25.64 $\uparrow$ +18.10 $\uparrow$ -2.54 $\downarrow$ +0.43 $\uparrow$ -1.29 $\downarrow$ +1.79 $\uparrow$
+CoT + Debias -8.84 $\downarrow$ -4.72 $\downarrow$ -3.71 $\downarrow$ -6.74 $\downarrow$ +24.03 $\uparrow$ +14.56 $\uparrow$ -4.35 $\downarrow$ -1.82 $\downarrow$ -0.60 $\downarrow$ +0.87 $\uparrow$
+HP + System 2 + CoT + Debias -9.19 $\downarrow$ -6.22 $\downarrow$ -5.29 $\downarrow$ -6.74 $\downarrow$ +25.64 $\uparrow$ +16.67 $\uparrow$ -5.72 $\downarrow$ -0.24 $\downarrow$ 0.00 +0.99 $\uparrow$
+Table 5: Results for GPT-4.
+Results reported compared to Standard Prompt (increased from the Standard prompt: $\uparrow$ in red, decrease: $\downarrow$ in green).
+Least stereotypical responses of each bias category are bolded.
+Avg.
+is the macro average of each prompting technique.
+B.3 Discussion of Mistral Results
+Table 6 shows Mistral-7B’s stereotypical engagement scores under various prompting techniques compared to a Standard Prompt baseline.
+Baseline and Overall Patterns.
+Under the Standard Prompt, Mistral-7B’s highest bias levels appear in Religion (53.23) and Beauty (56.82), with moderate values in domains like Profession (46.89) and Race (43.42).
+Although not as pronounced as with some larger models, these baselines provide a starting point to measure improvements.
+Influence of Human Persona.
+CoT alone tends to slightly increase bias averages, while System 1 and System 2 techniques deliver mixed outcomes.
+System 2, however, often yields more consistent bias reductions than System 1, improving categories like Beauty and Race.
+Adding a Human Persona (HP) further shapes the results—— HP combined with System 2, for example, significantly reduces bias in areas, including Beauty (-13.23↓) and Profession (-5.79↓).
+Effectiveness of Debiasing Techniques.
+Debiasing methods generally help decrease bias, especially when combined with System 2 and HP.
+Configurations like HP + System 2 + Debias and HP + System 2 + CoT + Debias stand out, offering broad improvements across multiple categories.
+The strongest reductions appear in areas that initially showed high bias, such as Beauty, Race, and Religion.
+Failure cases.
+A review of Table 6 highlights multiple instances where certain methods increase biases relative to the standard prompt.
+For example, CoT produces higher biases in Beauty, Beauty-P, Gender, Institution, Nationality, Profession, Race, and Religion; System 1 shows increase in bias for Beauty-P, Institution, Profession, Race, and Religion; and HP + System 1 amplifies biases in Beauty-P and Profession.
+HP + System 2 + CoT leads to greater bias in Profession and notably Religion, while MP + System 2 + CoT exhibits a particularly large increase in Religion.
+Standard Prompt + Debias also shows rises for Age, Beauty-P, Gender, Institution, Nationality, Profession, and Race, and CoT + Debias increases bias in Beauty-P, Gender, Institution, Nationality, Profession, and Race.
+In summary, Mistral-7B demonstrates that thoughtful prompt design—particularly leveraging System 2, Human Persona integration, and explicit debiasing—can substantially lower stereotypical responses across various domains.
+Type Age Beauty Beauty-P. Gender Insti. Nation. Prof. Race Religion Avg.
+Standard Prompt 32.00 56.82 38.02 50.30 38.59 43.23 46.89 43.42 53.23 44.72
+CoT -0.92 $\downarrow$ +0.09 $\uparrow$ +1.48 $\uparrow$ +1.42 $\uparrow$ +1.73 $\uparrow$ +1.52 $\uparrow$ +7.97 $\uparrow$ +2.06 $\uparrow$ +3.13 $\uparrow$ +2.06 $\uparrow$
+System 1 -1.52 $\downarrow$ -4.07 $\downarrow$ +0.24 $\uparrow$ -6.58 $\downarrow$ +1.05 $\uparrow$ -1.31 $\downarrow$ +0.91 $\uparrow$ +3.30 $\uparrow$ +3.37 $\uparrow$ -0.51 $\downarrow$
+System 2 -2.35 $\downarrow$ -5.91 $\downarrow$ +0.40 $\uparrow$ +0.75 $\uparrow$ +0.85 $\uparrow$ -0.75 $\downarrow$ -1.08 $\downarrow$ -1.38 $\downarrow$ -5.77 $\downarrow$ -1.69 $\downarrow$
+HP + System 1 -0.72 $\downarrow$ -6.95 $\downarrow$ +1.35 $\uparrow$ -5.67 $\downarrow$ -0.37 $\downarrow$ -2.66 $\downarrow$ +5.72 $\uparrow$ -3.64 $\downarrow$ -3.23 $\downarrow$ -1.95 $\downarrow$
+HP + System 2 -0.80 $\downarrow$ -13.23 $\downarrow$ -1.97 $\downarrow$ -3.44 $\downarrow$ -2.70 $\downarrow$ -4.70 $\downarrow$ -5.79 $\downarrow$ -4.97 $\downarrow$ +0.77 $\uparrow$ -4.09 $\downarrow$
+HP + CoT -2.87 $\downarrow$ -4.12 $\downarrow$ +0.06 $\uparrow$ -4.85 $\downarrow$ +0.10 $\uparrow$ -1.61 $\downarrow$ -2.39 $\downarrow$ +0.85 $\uparrow$ -4.66 $\downarrow$ -2.16 $\downarrow$
+HP + System 2 + CoT -3.08 $\downarrow$ -12.00 $\downarrow$ -2.96 $\downarrow$ -0.66 $\downarrow$ -2.40 $\downarrow$ -5.35 $\downarrow$ +0.26 $\uparrow$ -8.86 $\downarrow$ +5.47 $\uparrow$ -3.29 $\downarrow$
+MP + System 1 -1.50 $\downarrow$ -2.35 $\downarrow$ +1.23 $\uparrow$ -5.99 $\downarrow$ +1.24 $\uparrow$ -1.24 $\downarrow$ -0.38 $\downarrow$ -1.87 $\downarrow$ -5.08 $\downarrow$ -1.77 $\downarrow$
+MP + System 2 -0.12 $\downarrow$ -9.14 $\downarrow$ -2.51 $\downarrow$ -5.43 $\downarrow$ -2.52 $\downarrow$ -2.64 $\downarrow$ -1.98 $\downarrow$ -5.03 $\downarrow$ -2.12 $\downarrow$ -3.50 $\downarrow$
+MP + CoT -0.80 $\downarrow$ -2.05 $\downarrow$ -0.08 $\downarrow$ -6.27 $\downarrow$ +0.50 $\uparrow$ -0.13 $\downarrow$ -1.21 $\downarrow$ +1.60 $\uparrow$ -9.11 $\downarrow$ -1.95 $\downarrow$
+MP + System 2 + CoT -3.79 $\downarrow$ -7.89 $\downarrow$ -1.26 $\downarrow$ -5.78 $\downarrow$ -0.50 $\downarrow$ -1.77 $\downarrow$ -1.11 $\downarrow$ -2.87 $\downarrow$ +10.41 $\uparrow$ -1.62 $\downarrow$
+Standard Prompt + Debias +0.82 $\uparrow$ -2.29 $\downarrow$ +0.54 $\uparrow$ +2.21 $\uparrow$ +1.73 $\uparrow$ +0.86 $\uparrow$ +3.48 $\uparrow$ +0.78 $\uparrow$ -7.08 $\downarrow$ +0.12 $\uparrow$
+HP + Debias -0.47 $\downarrow$ -12.32 $\downarrow$ -0.23 $\downarrow$ -2.79 $\downarrow$ +0.51 $\uparrow$ -1.71 $\downarrow$ -3.36 $\downarrow$ -8.00 $\downarrow$ -5.23 $\downarrow$ -3.73 $\downarrow$
+System 2 + Debias -1.73 $\downarrow$ -7.90 $\downarrow$ +0.90 $\uparrow$ -3.96 $\downarrow$ +0.25 $\uparrow$ -1.39 $\downarrow$ -0.91 $\downarrow$ -7.53 $\downarrow$ -4.11 $\downarrow$ -2.93 $\downarrow$
+HP + System 2 + Debias -1.09 $\downarrow$ -15.75 $\downarrow$ -0.03 $\downarrow$ -7.07 $\downarrow$ -2.61 $\downarrow$ -2.92 $\downarrow$ -2.32 $\downarrow$ -6.75 $\downarrow$ +1.32 $\uparrow$ -3.98 $\downarrow$
+CoT + Debias -1.69 $\downarrow$ -2.49 $\downarrow$ +0.38 $\uparrow$ +1.13 $\uparrow$ +3.09 $\uparrow$ +0.63 $\uparrow$ +2.00 $\uparrow$ +2.27 $\uparrow$ -5.15 $\downarrow$ +0.02 $\uparrow$
+HP + System 2 + CoT + Debias -2.41 $\downarrow$ -12.06 $\downarrow$ -3.16 $\downarrow$ -3.68 $\downarrow$ -1.05 $\downarrow$ -4.35 $\downarrow$ -2.68 $\downarrow$ -8.44 $\downarrow$ -8.79 $\downarrow$ -5.18 $\downarrow$
+Table 6: Results for Mistral-7B.
+Results reported compared to Standard Prompt (increased from the Standard prompt: $\uparrow$ in red, decrease: $\downarrow$ in green).
+Least stereotypical responses of each bias category are bolded.
+Avg.
+is the macro average of each prompting technique.
+B.4 Discussion of Gemma3 Results
+Table 7 summarizes how Gemma3-27B’s stereotypical biases shift under various prompting and debiasing strategies when compared to a Standard Prompt baseline.
+We observe notable baseline biases in areas like Beauty (72.45) and Gender (69.46), while other categories like Institutional (27.59) and Nationality (40.50) start lower.
+Adjusting Prompting Strategies.
+Introducing System 1 and System 2, or adding CoT, produces mixed results.
+For example, System 1 slightly reduces overall bias on average (-1.07↓), and System 2 also shows moderate overall improvements (-0.78↓).
+Combining a Human Persona (HP) with System 2 is particularly effective, offering broad and notable reductions across multiple categories, including substantial drops in Age (-7.12↓), Beauty (-3.82↓), and Profession (-4.88↓).
+Impact of Debiasing.
+Debiasing methods generally lead to meaningful gains.
+Techniques like HP + System 2 + Debias achieve stronger reductions across many bias categories, notably Beauty (-6.77↓) and Race (-11.06↓).
+Adding CoT to System 2 and debiasing strategies typically enhances these improvements.
+Failure cases.
+Table 7 shows that several techniques inadvertently produce higher bias in certain categories compared to the standard prompt.
+CoT exhibits increases for Age, Beauty, Beauty-P., Gender, Institution, and Profession.
+System 1 raises bias in Beauty, Gender, and Profession, while System 2 inflates Beauty-P., Profession, and Religion.
+HP + System 1 amplifies Beauty, Gender, Profession, and Religion, and HP + CoT increases Beauty-P., Gender, and Profession.
+MP + System 1 similarly adds bias in Beauty, Gender, and Profession, MP + System 2 shows gains in Beauty and Profession, and MP + CoT amplifies Beauty, Gender, Profession, and Religion.
+MP + System 2 + CoT raises biases in Beauty, Profession, and Religion, and Standard Prompt + Debias exhibits increases for Institution and Religion.
+Even in some debias configurations, such as HP + Debias, System 2 + Debias, and CoT + Debias, the techniques inflate Beauty-P.
+or Profession.
+Type Age Beauty Beauty-P. Gender Insti. Nation. Prof. Race Religion Avg.
+Standard Prompt 42.06 72.45 48.08 69.46 27.59 40.50 63.72 60.73 57.14 53.53
+CoT +0.03 $\uparrow$ +3.48 $\uparrow$ +0.64 $\uparrow$ +7.18 $\uparrow$ +0.10 $\uparrow$ -1.90 $\downarrow$ +4.85 $\uparrow$ -0.25 $\downarrow$ -2.21 $\downarrow$ +1.32 $\uparrow$
+System 1 -0.25 $\downarrow$ +1.93 $\uparrow$ -0.89 $\downarrow$ +1.25 $\uparrow$ -1.15 $\downarrow$ -3.53 $\downarrow$ +2.28 $\uparrow$ -5.48 $\downarrow$ -3.72 $\downarrow$ -1.07 $\downarrow$
+System 2 -2.43 $\downarrow$ -0.36 $\downarrow$ +1.00 $\uparrow$ -2.66 $\downarrow$ -0.67 $\downarrow$ -2.93 $\downarrow$ +0.79 $\uparrow$ -6.49 $\downarrow$ +6.75 $\uparrow$ -0.78 $\downarrow$
+HP + System 1 -3.44 $\downarrow$ +1.69 $\uparrow$ -1.21 $\downarrow$ +1.08 $\uparrow$ -0.92 $\downarrow$ -6.28 $\downarrow$ +4.14 $\uparrow$ -7.12 $\downarrow$ +2.01 $\uparrow$ -1.12 $\downarrow$
+HP + System 2 -7.12 $\downarrow$ -3.82 $\downarrow$ -1.07 $\downarrow$ -6.65 $\downarrow$ -2.22 $\downarrow$ -6.67 $\downarrow$ -4.88 $\downarrow$ -6.64 $\downarrow$ -0.98 $\downarrow$ -3.62 $\downarrow$
+HP + CoT -3.65 $\downarrow$ -0.12 $\downarrow$ +0.65 $\uparrow$ +2.31 $\uparrow$ -0.81 $\downarrow$ -4.93 $\downarrow$ +1.29 $\uparrow$ -6.20 $\downarrow$ -7.84 $\downarrow$ -2.15 $\downarrow$
+HP + System 2 + CoT -2.49 $\downarrow$ -3.89 $\downarrow$ -1.52 $\downarrow$ -3.91 $\downarrow$ -1.99 $\downarrow$ -5.86 $\downarrow$ -1.30 $\downarrow$ -9.96 $\downarrow$ -2.85 $\downarrow$ -3.76 $\downarrow$
+MP + System 1 -4.35 $\downarrow$ +4.19 $\uparrow$ -2.29 $\downarrow$ +2.25 $\uparrow$ -0.93 $\downarrow$ -4.30 $\downarrow$ +3.03 $\uparrow$ -4.29 $\downarrow$ -1.73 $\downarrow$ -1.25 $\downarrow$
+MP + System 2 -4.36 $\downarrow$ +0.08 $\uparrow$ -2.81 $\downarrow$ -2.53 $\downarrow$ -0.90 $\downarrow$ -5.50 $\downarrow$ +1.25 $\uparrow$ -7.11 $\downarrow$ -1.88 $\downarrow$ -2.64 $\downarrow$
+MP + CoT -4.77 $\downarrow$ +3.25 $\uparrow$ -2.55 $\downarrow$ +2.03 $\uparrow$ -1.05 $\downarrow$ -5.51 $\downarrow$ +5.28 $\uparrow$ -4.86 $\downarrow$ +3.39 $\uparrow$ -0.54 $\downarrow$
+MP + System 2 + CoT -5.41 $\downarrow$ +3.23 $\uparrow$ -4.47 $\downarrow$ -0.11 $\downarrow$ -1.39 $\downarrow$ -5.57 $\downarrow$ +2.09 $\uparrow$ -7.99 $\downarrow$ +2.01 $\uparrow$ -1.96 $\downarrow$
+Standard Prompt + Debias -2.46 $\downarrow$ -3.53 $\downarrow$ -2.84 $\downarrow$ -3.08 $\downarrow$ +0.18 $\uparrow$ -4.22 $\downarrow$ -0.11 $\downarrow$ -7.22 $\downarrow$ +1.88 $\uparrow$ -2.91 $\downarrow$
+HP + Debias -1.10 $\downarrow$ -6.37 $\downarrow$ +1.15 $\uparrow$ -0.86 $\downarrow$ -1.06 $\downarrow$ -1.03 $\downarrow$ -2.10 $\downarrow$ -12.59 $\downarrow$ -2.97 $\downarrow$ -3.00 $\downarrow$
+System 2 + Debias -1.29 $\downarrow$ -1.71 $\downarrow$ +1.96 $\uparrow$ -4.33 $\downarrow$ -1.28 $\downarrow$ -2.58 $\downarrow$ +0.10 $\uparrow$ -10.62 $\downarrow$ -3.72 $\downarrow$ -2.50 $\downarrow$
+HP + System 2 + Debias -4.33 $\downarrow$ -6.77 $\downarrow$ -2.43 $\downarrow$ -3.07 $\downarrow$ -2.04 $\downarrow$ -3.78 $\downarrow$ -1.88 $\downarrow$ -11.06 $\downarrow$ -0.20 $\downarrow$ -3.96 $\downarrow$
+CoT + Debias -3.59 $\downarrow$ -3.44 $\downarrow$ +0.41 $\uparrow$ -3.77 $\downarrow$ -0.50 $\downarrow$ -1.33 $\downarrow$ -2.53 $\downarrow$ -12.36 $\downarrow$ -5.03 $\downarrow$ -3.57 $\downarrow$
+HP + System 2 + CoT + Debias -2.26 $\downarrow$ -6.38 $\downarrow$ -1.00 $\downarrow$ -2.66 $\downarrow$ -1.74 $\downarrow$ -2.66 $\downarrow$ -0.44 $\downarrow$ -11.68 $\downarrow$ -3.72 $\downarrow$ -3.62 $\downarrow$
+Table 7: Results for Gemma3-27B.
+Results reported compared to Standard Prompt (increased from the Standard prompt: $\uparrow$ in red, decrease: $\downarrow$ in green).
+Least stereotypical responses of each bias category are bolded.
+Avg.
+is the macro average of each prompting technique.
+B.5 Discussion of Llama3.3 Results
+Table 8 shows how Llama3.3-70B’s stereotypical biases shift under a range of prompting and debiasing techniques, using the Standard Prompt as a baseline.
+Initially, categories like Beauty (75.15) and Gender (74.68) display relatively high bias scores, while others such as Age (40.85) and Institutional (36.57) begin at more moderate levels.
+Influence of Human Persona.
+Introducing CoT or System 1 or 2 reduces biases for most categories.
+System 2, for example, cuts biases notably across various domains (Avg.
+-4.38↓), while CoT alone also lowers overall bias (-3.65↓).
+Incorporating a Human Persona (HP) further amplifies these improvements.
+HP combined with System 2 consistently yields stronger reductions, bringing down biases in challenging domains like Beauty (-22.49↓) and Profession (-12.29↓).
+Adding CoT to HP + System 2 often enhances these positive effects.
+Effectiveness of Debiasing.
+Debiasing techniques substantially reduce stereotypical engagement, especially when combined with other strategies.
+Configurations like HP + System 2 + Debias or HP + System 2 + CoT + Debias yield large, broad-based reductions.
+For instance, HP + System 2 + CoT + Debias reduces Beauty (-32.81↓), Profession (-17.82↓), and Religion (-20.94↓) biases, making it one of the most effective configurations tested.
+Failure cases.
+Table 8 shows a few specific cases where biases increase relative to the standard prompt.
+CoT inflates Institution bias by 2.62, while System 1 raises bias in Beauty-P.
+(1.03), Profession (0.84), and Race (0.15).
+Standard Prompt + Debias also shows a small increase in Institution (0.31).
+In summary, Llama3.3-70B shows less stereotypical engagement when integrated approaches that combine human persona, and explicit debiasing methods.
+These strategies substantially mitigate stereotypical responses across various bias categories, indicating that tailored prompt engineering can help achieve more balanced and less biased model outputs.
+Type Age Beauty Beauty-P. Gender Insti. Nation. Prof. Race Religion Avg.
+Standard Prompt 40.85 75.15 45.21 74.68 36.57 47.37 67.89 58.69 50.79 55.24
+CoT -2.44 $\downarrow$ -6.94 $\downarrow$ -3.54 $\downarrow$ -6.54 $\downarrow$ +2.62 $\uparrow$ -3.62 $\downarrow$ -1.46 $\downarrow$ -4.28 $\downarrow$ -6.72 $\downarrow$ -3.65 $\downarrow$
+System 1 -2.96 $\downarrow$ -1.08 $\downarrow$ +1.03 $\uparrow$ -0.10 $\downarrow$ -0.61 $\downarrow$ -2.99 $\downarrow$ +0.84 $\uparrow$ +0.15 $\uparrow$ -11.98 $\downarrow$ -1.96 $\downarrow$
+System 2 -3.23 $\downarrow$ -6.71 $\downarrow$ -3.66 $\downarrow$ -5.48 $\downarrow$ -1.72 $\downarrow$ -3.25 $\downarrow$ -3.30 $\downarrow$ -3.52 $\downarrow$ -8.60 $\downarrow$ -4.38 $\downarrow$
+HP + System 1 -4.09 $\downarrow$ -15.87 $\downarrow$ -11.28 $\downarrow$ -6.90 $\downarrow$ -3.37 $\downarrow$ -7.13 $\downarrow$ -7.36 $\downarrow$ -7.43 $\downarrow$ -16.41 $\downarrow$ -8.87 $\downarrow$
+HP + System 2 -8.40 $\downarrow$ -22.49 $\downarrow$ -20.58 $\downarrow$ -10.10 $\downarrow$ -2.46 $\downarrow$ -5.33 $\downarrow$ -12.29 $\downarrow$ -10.32 $\downarrow$ -11.40 $\downarrow$ -11.48 $\downarrow$
+HP + CoT -7.71 $\downarrow$ -19.25 $\downarrow$ -17.46 $\downarrow$ -12.65 $\downarrow$ -3.94 $\downarrow$ -6.52 $\downarrow$ -9.02 $\downarrow$ -6.42 $\downarrow$ -11.73 $\downarrow$ -10.52 $\downarrow$
+HP + System 2 + CoT -10.08 $\downarrow$ -24.51 $\downarrow$ -22.19 $\downarrow$ -15.68 $\downarrow$ -3.93 $\downarrow$ -9.11 $\downarrow$ -14.02 $\downarrow$ -13.85 $\downarrow$ -7.04 $\downarrow$ -13.37 $\downarrow$
+MP + System 1 -5.13 $\downarrow$ -12.78 $\downarrow$ -11.68 $\downarrow$ -12.76 $\downarrow$ -3.47 $\downarrow$ -4.08 $\downarrow$ -6.79 $\downarrow$ -5.69 $\downarrow$ -13.87 $\downarrow$ -8.47 $\downarrow$
+MP + System 2 -7.52 $\downarrow$ -25.49 $\downarrow$ -20.95 $\downarrow$ -7.19 $\downarrow$ -6.90 $\downarrow$ -9.84 $\downarrow$ -14.59 $\downarrow$ -9.49 $\downarrow$ -11.11 $\downarrow$ -12.45 $\downarrow$
+MP + CoT -6.55 $\downarrow$ -18.70 $\downarrow$ -16.50 $\downarrow$ -9.31 $\downarrow$ -4.11 $\downarrow$ -5.33 $\downarrow$ -9.78 $\downarrow$ -7.71 $\downarrow$ -13.09 $\downarrow$ -10.12 $\downarrow$
+MP + System 2 + CoT -9.21 $\downarrow$ -24.01 $\downarrow$ -21.22 $\downarrow$ -11.43 $\downarrow$ -6.61 $\downarrow$ -8.74 $\downarrow$ -14.69 $\downarrow$ -10.75 $\downarrow$ -11.81 $\downarrow$ -13.16 $\downarrow$
+Standard Prompt + Debias -3.03 $\downarrow$ -6.36 $\downarrow$ -2.37 $\downarrow$ -2.94 $\downarrow$ +0.31 $\uparrow$ -2.34 $\downarrow$ -2.58 $\downarrow$ -4.38 $\downarrow$ -10.17 $\downarrow$ -3.76 $\downarrow$
+HP + Debias -11.10 $\downarrow$ -28.40 $\downarrow$ -15.53 $\downarrow$ -12.65 $\downarrow$ -5.14 $\downarrow$ -8.55 $\downarrow$ -16.78 $\downarrow$ -19.73 $\downarrow$ -20.36 $\downarrow$ -15.24 $\downarrow$
+System 2 + Debias -8.03 $\downarrow$ -24.77 $\downarrow$ -15.70 $\downarrow$ -14.01 $\downarrow$ -2.51 $\downarrow$ -4.65 $\downarrow$ -12.13 $\downarrow$ -13.83 $\downarrow$ -13.48 $\downarrow$ -12.12 $\downarrow$
+HP + System 2 + Debias -10.12 $\downarrow$ -32.06 $\downarrow$ -22.62 $\downarrow$ -19.64 $\downarrow$ -4.85 $\downarrow$ -11.78 $\downarrow$ -19.04 $\downarrow$ -18.88 $\downarrow$ -15.50 $\downarrow$ -17.16 $\downarrow$
+CoT + Debias -8.77 $\downarrow$ -15.43 $\downarrow$ -14.47 $\downarrow$ -9.72 $\downarrow$ -2.84 $\downarrow$ -6.33 $\downarrow$ -10.08 $\downarrow$ -15.34 $\downarrow$ -20.63 $\downarrow$ -11.51 $\downarrow$
+HP + System 2 + CoT + Debias -12.27 $\downarrow$ -32.81 $\downarrow$ -24.43 $\downarrow$ -19.26 $\downarrow$ -5.57 $\downarrow$ -12.40 $\downarrow$ -17.82 $\downarrow$ -18.71 $\downarrow$ -20.94 $\downarrow$ -18.24 $\downarrow$
+Table 8: Results for Llama3.3-70B.
+Results reported compared to Standard Prompt (increased from the Standard prompt: $\uparrow$ in red, decrease: $\downarrow$ in green).
+Least stereotypical responses of each bias category are bolded.
+Avg.
+is the macro average of each prompting technique.
+Appendix C Statistical Test
+Group 1 Group 2 Model $\tau$ $p$ Bias Type
+HP Debias Standard Prompt GPT-4o-mini 0.175 <<0.001 Ageism
+HP+System 2+CoT+Debias Standard Prompt GPT-4o-mini 0.193 0.168 Ageism
+HP + Debias Standard Prompt GPT-4o-mini 0.126 <<0.001 Race
+HP + System 2 + Debias Standard Prompt GPT-4o-mini 0.075 <<0.001 Beauty Prof.
+HP Debias Standard Prompt Gemma3 0.376 <<0.001 Race
+HP + System 2 Standard Prompt Gemma3 0.111 0.092 Insti.
+HP + System 2 + Debias Standard Prompt Gemma3 0.216 <<0.001 Beauty
+HP + System 2 Standard Prompt Gemma3 0.184 <<0.001 Nation
+HP+System 2+CoT+Debias Standard Prompt Llama3.3 0.159 <<0.001 Religion
+HP + System 2 Standard Prompt Llama3.3 -0.276 0.290 Insti.
+HP + Debias Standard Prompt Llama3.3 0.456 <<0.001 Ageism
+HP + System 2 + Debias Standard Prompt Llama3.3 0.093 <<0.001 Beauty Prof.
+Table 9: Kendall’s $\tau$ test results where we try to see if group 1’s stereotypical engagement is less than of group 2 (Standard Prompt).
+We use a significance level of $\alpha<0.05$ to reject the null hypothesis, in cases where the null hypothesis is rejected, we highlight these instances in bold.
+Appendix D Does CoT Prompting Best Model System 2?
 Prompting Techniques $\tau$ $p$ $H_{0}$?
@@
 Now we further investigate whether CoT prompting is most similar to the way that LLMs model System 2 reasoning.
-While Figure 2 shows that the stereotyping rate of CoT is most similar to System 1 prompts, these may be from different test items.
+While Figure 2 shows that the stereotyping rate of CoT is most similar to System 1 and Standard prompts, these may be from different test items.
 Here we tackle this question directly by computing the Kendall $\tau$ coefficient (Kendall, 1938) between CoT-prompted responses and those of the other variants.
@@
 While prior work has found that CoT prompting leads to better multi-step mathematical and formal reasoning capabilities (Wei et al., 2022; Yu et al., 2023; Wang et al., 2023), that align with System 2 cognitive processes, the growing body of evidence suggests that this is because the formal reasoning setting contextualizes LLMs to generate text reflecting System 2 reasoning in people.
-7 Invalid LLMs Responses
+Appendix E Invalid LLMs Responses
 We excluded certain examples due to the language models providing invalid responses.
@@
 When calculating the prevalence of stereotypical responses, we consider these responses, which demonstrate awareness of stereotypes, as anti-stereotype responses.
-8 Conclusion
-Our study has contributed to the understanding and reduction of social biases in LLMs through prompting techniques grounded in dual process theory.
-By harnessing the cognitive frameworks of System 1 and System 2, as well as the incorporation of human-like personas and debiasing prompts, our research not only clarifies the role of these cognitive processes in LLMs but also demonstrates practical methods for reducing biases.
-Our findings reveal that System 2 prompts, particularly when combined with a Human Persona, consistently reduce stereotypical judgments across various social bias categories.
-Biases were further reduced when using a debiasing prompt, which can be seen as a social bias-focused System 2 prompt, along with a human persona.
-This indicates a profound potential for combining analytical thought processes and personalized prompting to enhance the ethical performance of LLMs.
-Furthermore, our use of different models and bias datasets has allowed us to explore the diverse applications of these techniques, ensuring our results are robust and applicable across different contexts.
-These findings underscore the potential of sophisticated prompting strategies in enhancing the ethical aspects of AI, pointing towards a future where LLMs can assist in creating more inclusive digital environments.
-Through continued exploration and refinement of these methods, we anticipate further advancements in the responsible deployment of AI technologies.